### v1

# Odd-One-Out: Anomaly Detection by Comparing with Neighbors

Abstract
This paper introduces a novel anomaly detection (AD) problem that focuses on identifying ‘odd-looking’ objects relative to the other instances within a scene. Unlike the traditional AD benchmarks, in our setting, anomalies in this context are scene-specific, defined by the regular instances that make up the majority. Since object instances are often partly visible from a single viewpoint, our setting provides multiple views of each scene as input. To provide a testbed for future research in this task, we introduce two benchmarks, ToysAD-8K and PartsAD-15K. We propose a novel method that generates 3D object-centric representations for each instance and detects the anomalous ones through a cross-examination between the instances. We rigorously analyze our method quantitatively and qualitatively in the presented benchmarks.
1 Introduction
Anomaly detection (AD) [13, 34] aims to detect patterns that deviate from expected behavior. In standard computer vision AD benchmarks, the non-conforming (or anomalous) patterns in images are often either due to high-level variations such as introduction of an object instance from an unseen category [2, 8, 12] or low-level variations in object shape and texture [6, 16] (see Fig. 1(a)). In these benchmarks, the definitions of normal and anomalous patterns are typically not predefined but implicitly learned through representations and/or classifiers [21, 2, 16] that are invariant to appearance changes within the normal data distribution while being sensitive to the anomalous ones. However, in many real-world applications such as quality control in production lines, product specification, hence the definition of ‘normality’, is not only known in advance but also specific to each object instance. For example, while a coffee cup with a red handle is considered normal and one with a blue handle is anomalous when the goal is to produce red-handled cups, but the opposite is true if the goal is to produce blue-handled cups. This instance-specific definition of normality cannot be addressed by methods that rely on a global and implicit definition of normality.
In this paper, inspired from real-world visual inspection scenarios, we introduce a new AD problem, along with two new benchmarks with different characteristics and a novel solution. As depicted in Fig. 1 (b), in this proposed setting, each image contains a scene with multiple instances of the same object (e.g., coffee cups) including either only normal, or a mix of normal and anomalous instances where the anomalous instances constitute the minority. As defining precise visual specifications for shape and appearance is often infeasible, we assume that normal instances, which form the majority, provide a scene-specific reference for ‘normality’. Our objective is to detect anomalous instances in a scene while generalizing to previously unseen scenes including novel objects and spatial configurations.
Unlike existing benchmarks, our task often requires a holistic understanding of the scene by comparing instances with each other, although some anomalies, such as cracks and misaligned parts, can be identified by inspecting individual instances. Additionally, since performing AD from a single view can be ambiguous due to self-occlusion and occlusion between the instances, we provide multiple views of the scene to cover its entire relevant extent, unlike the existing benchmarks that provide a single view as input (see Fig. 1(c)). Our goal is to detect anomalous samples in previously unseen scenes from multiple views.
Figure 1: The standard and our new AD settings.
The proposed problem presents several challenges requiring the following capabilities: i) 3D understanding of the scene and registering the views from multiple camera viewpoints without groundtruth 3D knowledge while identifying potential occlusions, ii) aligning and comparing object instances with each other without their relative pose information in both training and evaluation, iii) learning representations that generalize to unseen object instances during testing. To address these challenges, we propose a novel method that takes as input multiple views of the same scene, projects them into a 3D voxel grid, produces a 3D object-centric representation for each instance, and predicts their labels by cross-correlating instances through an efficient attention mechanism. We leverage recent advances in differentiable rendering [31] and self-supervised learning [33] to supervise 3D representation learning. Specifically, we render the voxel representations for multiple viewpoints and match them with the given 2D views to produce geometrically consistent features. Additionally, we encourage our model to learn part-aware 3D representations by distilling features from 2D self-supervised model DINOv2 [33], enhancing the correspondence matching across instances. Finally, since there is no prior benchmark for this task, we propose two new benchmarks – ToysAD-8K and PartsAD-15K, including instances from common semantic object categories as well as mechanical parts respectively to provide a testbed for future research. Our method significantly outperforms various baselines that do not compare instances with each other, and we rigorously analyze various aspects of our model and benchmarks.
2 Related Work
AD benchmarks. A key challenge in AD research is the scarcity of large datasets containing realistic anomalies. Earlier works [11, 36] focusing on high-level semantic anomalies often use existing classification datasets by treating a subset of classes as anomalies and the remainder as normal. There also exist several datasets containing real-world anomaly instances. For example, MVTec-AD [5] includes industrial objects with various defects like scratches, dents, and contaminations, Carrera et al. missing[10] presents various defects in nanofibrous material, and VisA [59] comprises complex industrial objects such as PCBs, as well as simpler objects like capsules and cashews, spanning a total of 12 categories. These datasets assume that objects are pose-aligned, and both normal images and their anomaly counterparts have the same pose. Zhou et al. missing[58] propose a pose-agnostic framework by introducing the PAD dataset, which comprises images of 20 LEGO bricks of animal toys from diverse viewpoints/poses. Unlike the works discussed above, we focus on AD in multi-object multi-view scene environments, where anomalies are predicted by assessing their mutual similarity with other objects in the scene. The proposed setting enables our framework to work seamlessly on novel object instances without requiring further training.
Few-shot AD. There are some works [18, 23, 52, 50, 7] aiming to detect anomalies from a small number of normal samples as support images. In our setting, the concept of normality in each image is also learned from a few instances only. However, unlike them, the concept of normality is scene-specific, and our method can generalize to previously unseen instances without requiring any modification by learning from a support set. In addition, our setting involves multi-object multi-view data samples as input, unlike the single-object single-view in theirs.
Multi-view 3D vision. Multi-view 3D detection [37, 40, 51, 48] is a related problem that aims to predict the locations and classes of objects in 3D space, given multi-view images of a scene along with their corresponding camera poses as input. Most existing works first project 2D image features onto a 3D voxel grid, followed by a detection head [29, 53, 9] that outputs the final 3D bounding boxes and class labels. While our task can be naively solved by treating it as a 3D object detection problem with anomaly and normal as the two possible classes, this approach has limited ability to perform effective comparisons with other instances in the scene, which is required for fine-grained AD. We compare our method to a 3D object detection technique in Sec. 4. Another related area involves taking multi-view images as input and training a feedforward model for 3D volume-based reconstruction [32, 55, 42] and novel view synthesis [45, 47, 54, 15, 24]. Similarly, in this work, we focus on learning a feedforward model in a multi-object scene environment using sparse multi-view images but for AD.
Leveraging foundation models. Large-scale pretraining on image datasets has shown impressive generalization capabilities on various tasks [35, 25, 33]. Previous works [57, 3] demonstrate that features extracted from DINOv2 [33] serve as effective dense visual descriptors with localized semantic information for dense correspondence estimation task. Some recent works [26, 46] use feature distillation techniques to leverage 2D foundation vision models for 3D task. Inspired by these works we utilize DINOv2 to distill its dense semantic knowledge into our 3D network that enables our network to infer robust local correspondences, which aids in fine-grained object matching.
3 Method
3.1 Overview
Consider a scene containing multiple rigid objects $\{o_{n}\}_{n=1}^{N}$ of the same instance, where $N$ represents the number of objects in the scene. Each object’s pose is arbitrary and unknown. The goal is to identify anomalies in the group of objects by assessing their mutual similarity. An observation in our setting consists of $M$-view $H\times W$ dimensional RGB images =$\mathcal{I}=\{\bm{I}_{t}\}_{t=1}^{M}$, and their corresponding camera projection matrices $\mathcal{P}=\{\bm{P}_{t}\}_{t=1}^{M}$, $\bm{P}_{t}=\bm{K}\left[\bm{R}_{t}|\bm{T}_{t}\right]$, with intrinsic $\bm{K_{t}}$, rotation $\bm{R}_{t}$ and translation $\bm{T}_{t}$ matrices. In our setting, $M$ is $5$, forming sparse-view inputs. Our goal is to learn a mapping $\psi$ from the multi-view images to object-centric anomaly labels ${y}_{n}\in\{0,1\}$ and its corresponding 3D bounding box $\bm{b}_{n}$, defined as:
$\psi:\{(\bm{I}_{t},\bm{P}_{t})\}_{t=1}^{M}\mapsto\{({y}_{n},\bm{b}_{n})\}_{n=1}^{N}.$ (1)
Note, labels $y_{n}$ are defined relative to other objects in the scene. For example, consider a group of three coffee cups, of which two have red handles, and one has a blue handle. In this example, the latter cup is considered an anomaly.
Figure 2: Overview of our framework. We extract features from a sequence of input views using a 2D CNN and then back-project them into a 3D volume, which is then refined using a 3D CNN, resulting in $\bm{F}_{v}$. We then extract object-centric feature volumes $\{\bm{z}_{n}\}_{n=1}^{N}$ using RoI pooling, where $\mathcal{B}$ estimates the RoI region of each object. The cross-instance matching module $\mathcal{M}$ then learns a correlation among the objects using a sparse voxel attention. To improve the 3D representation of the scene, we distill the knowledge of a 2D vision model namely DINOv2, and integrate the learned knowledge into our 3D network via differentiable rendering.
Our architecture, illustrated in Fig. 2, is comprised of three main components: First, the 3D feature fusion module encodes each view image and projects it to 3D, forming a fused 3D feature volume. Second, the feature distillation block is employed to enhance the 3D feature volume through differentiable rendering. This facilitates finding local correspondences between objects. Finally, the cross-instance matching module leverages established correspondences to compare all similar object regions in the scene using a sparse voxel attention mechanism. Next, we elaborate on details.
3.2 3D Feature Volume Construction
We first extract 2D features $\bm{F}_{t}=\mathcal{E}_{2D}(\bm{I}_{t})\in\mathbb{R}^{d\times h\times w}$ for each input view using a shared CNN encoder $\mathcal{E}_{2D}$, where $d$ is the feature dimension. These 2D features are then projected into 3D voxel space as follows:
$\bm{F}_{v}=\mathcal{E}_{3D}(\texttt{aggr}(\{\Pi_{\texttt{proj}}(\bm{F}_{t},\bm{P}_{t})\}_{t=1}^{M})),$ (2)
where $\Pi_{\texttt{proj}}$ back-projects each view feature $\bm{F}_{t}$ using known camera intrinsic and extrinsic, generating 3D feature volumes of size $d\times v_{x}\times v_{y}\times v_{z}$. These feature volumes are aggregated over all input views using an average operation as in [32, 42]. Finally, a 3D CNN-based network $\mathcal{E}_{3D}$ is employed to refine the aggregated feature volume, resulting in a final voxel representation $\bm{F}_{v}$.
We use volume rendering [31] to reconstruct the geometry and appearance of the scene. We implement the rendering operation as in [24], where for each 3D query point on a ray, we retrieve its corresponding features by bilinearly interpolating between the neighboring voxel grids. Specifically, we first apply a two-layered $1\times 1\times 1$ convolution block, i.e., $\alpha_{c}$ and $\alpha_{\sigma}$ respectively, to obtain color and density volumes denoted as ($\bm{V}_{c}$, $\bm{V}_{\sigma}$). Then, the pixel-wise color and density maps are composed by integrating along a camera ray using volume renderer $\mathcal{R}$. Following this, we compute the L2 image reconstruction loss $\mathcal{L}^{\text{im}}_{t}$. Formally, the image rendering and its corresponding loss function for a single viewpoint $\bm{P}_{t}$ are shown below:
$[\hat{\bm{I}}_{t},\hat{\bm{I}}_{t\sigma}]=\mathcal{R}([\bm{V}_{c},\bm{V}_{\sigma}],\bm{P}_{t}),\quad\mathcal{L}^{\text{im}}_{t}=||\bm{I}_{t}-\hat{\bm{I}}_{t}||^{2}+\lambda_{\sigma}||\bm{I}_{t\sigma}-\hat{\bm{I}}_{t\sigma}||^{2},$ (3)
where $\hat{\bm{I}}_{t}$ and $\hat{\bm{I}}_{t\sigma}$ are rendered image and mask. $\lambda_{\sigma}$ is a loss weight.
We propose to improve the 3D voxel representation $\bm{F}_{v}$ by also reconstructing neural features instead of just color and density. We supervise the feature reconstruction by a pretrained 2D image encoder $\Phi$ as a teacher network. We choose DINOv2 [33] as the teacher network due to its excellent ability to capture various object geometries and correspondences.
We use a projector function $\beta$, implemented as a four-layered $1\times 1\times 1$ convolution block that projects $\bm{F}_{v}$ to a neural feature field $\bm{V}_{f}$, changing the channel dimension from $d$ to $d_{f}$. Similar to color rendering, we generate rendered features $\hat{\Phi}_{t}$ of size $d_{f}\times h_{f}\times w_{f}$ at a given viewpoint using volume renderer $\mathcal{R}$. The objective is to minimize the difference between the rendered features and the teacher’s features $\Phi(\bm{I}_{t})$. We choose cosine distance as our feature loss ($\mathcal{L}^{\text{feat}}_{t}$) which we find easier to optimize compared to the standard L2 loss. We apply stop-gradient to density in rendering of features $\hat{\Phi}_{t}$, as the teacher’s features are not fully multi-view consistent [3], which could harm the quality of reconstructed geometry. The final reconstruction loss is the sum of all image and feature reconstruction losses weighed by a factor $\lambda_{f}$:
$\mathcal{L}^{r}=\sum_{t=1}^{M}(\mathcal{L}^{\text{im}}_{t}+\lambda_{f}\mathcal{L}^{\text{feat}}_{t}).$ (4)
The key benefits of reconstructing DINOv2 features are twofold. 1) Firstly, distilling features from general-purpose feature extractors pre-trained on large external datasets, incorporates open-world knowledge into the 3D representation. This enables our model to perform significantly better on unseen object instances or even on novel categories, as demonstrated in the experiments. 2) Secondly, the distillation enforces consistent 3D scene representation, leading to identical features for the same object geometries. This enables the model to infer robust local correspondences (see Fig. 4), which aids in fine-grained object matching.
3.3 Object-centric 3D Features Extraction
We extract bounding box regions of objects using a box estimator $\mathcal{B}$. First, we obtain a voxel reconstruction of the scene by applying a threshold to the predicted density $\bm{V}_{\sigma}$. Then, we employ DBScan [20], a density-based clustering method to retrieve all bounding box regions $\{\bm{b}_{n}\}_{n=1}^{N}$ corresponding to the objects in the scene. Next we obtain object-centric feature volumes $\{\bm{z}_{n}\}_{n=1}^{N}$, each with a size of $8\times 8\times 8$ from these regions through applying RoI pooling [17].
To extract correspondence indices given two objects with feature volumes $\bm{z}_{n}$ and $\bm{z}_{m}$, We formally define a function $\mathcal{C}_{k}$:
$\mathcal{C}_{k}(\bm{z}_{n},\bm{z}_{m})=\texttt{top}_{k}[\beta(\bm{z}_{n})^{T}\beta(\bm{z}_{m})].$ (5)
The function returns top-$k$ most relevant feature locations in $\bm{z}_{m}$ for each voxel location in $\bm{z}_{n}$. This is achieved by first projecting each voxel feature using the same projector function $\beta$ and then computing their voxel-level pairwise similarity. We use $\mathcal{C}_{k}$ to perform sparse attention-based comparisons between multiple object volumes, as described next.
3.4 Cross-instance Matching
The cross-instance matching module $\mathcal{M}$ takes the object features $\{\bm{z}_{n}\}_{n=1}^{N}$, effectively learns to correlate them using sparse voxel attention, and subsequently predicts their object-specific labels.
Unlike the vanilla self-attention module in standard transformers [19] that uses all tokens for the attention computation, which is inefficient for our task and may introduce noisy interactions with irrelevant features, potentially degrading performance. To overcome this, we compute the sparse voxel attention only among geometrically corresponding voxel locations, as shown in the inset.
Let $z_{n}[i]\in\mathbb{R}^{d}$ denote the $i$-th voxel of the $n$-th object volume in the scene. The query, key, and value embeddings are calculated using linear projections as:
$\bm{Q}_{n}[i]=\bm{W}^{Q}{\bm{z}}_{n}[i],\quad\bm{K}_{n}[i]=\bm{W}^{K}{\bm{z}}_{n}[i],\quad\bm{V}_{n}[i]=\bm{W}^{V}{\bm{z}}_{n}[i],$ (6)
where the weights $\bm{W}^{Q}$, $\bm{W}^{K}$ and $\bm{W}^{V}$ are shared across all objects.
Let $c_{k}^{nm}[i]$ denote the set of all corresponding voxel indices obtained using Eq. 5. Then, the attention is calculated as:
$\bar{\bm{z}}_{n}[i]=\sum_{\begin{subarray}{c}m=1\\ m\neq n\end{subarray}}^{N}\sum_{j\in c_{k}^{nm}[i]}\text{softmax}\left(\frac{\bm{Q}_{n}[i]\bm{K}_{m}[j]}{\sqrt{d}}\right)\bm{V}_{m}[j].$ (7)
The updated feature volume $\bar{\bm{z}}_{n}$ is passed through 3D CNN blocks to downsample by a factor of $1/8$, which is finally reshaped into a vector and fed to a 2-layer MLP outputting the final prediction $\hat{y}_{n}$. The classification loss is calculated as:
$\mathcal{L}^{\text{bce}}=\sum_{n=1}^{N}\ell_{\text{bce}}(\hat{y}_{n},y_{n}),$ (8)
where $\ell_{\text{bce}}$ is the binary cross-entropy loss function. The total training loss of our framework is $\mathcal{L}=\mathcal{L}^{\text{bce}}+\lambda_{p}\mathcal{L}^{r}$, where $\lambda_{p}$ is a loss weight. We employ stage-wise training: first pretraining with only the reconstruction loss ($\mathcal{L}^{r}$), followed by end-to-end training with both losses.
4 Experiments
4.1 Datasets
As no prior dataset exists for the task, we propose two challenging scene datasets, both essential for thorough evaluation: ToysAD-8K and PartsAD-15K. ToysAD-8K includes real-world objects from multiple categories. This allows us to evaluate our model’s ability to generalize to unseen object categories. PartsAD-15K comprises a more diverse collection of mechanical object parts with arbitrary shapes, thus being free from any class-level inductive biases. Both datasets include a wide range of fine-grained anomaly instances motivated by real-world applications in inspection and quality control. Scenes are generated with diverse backgrounds, illuminations, and camera viewpoints using photo-realistic ray tracing [43]. Next, we discuss the data generation for both datasets.
ToysAD-8K. We begin with a subset of $1050$ shapes from the Toys4K dataset [41]. The subset covers a wide range of objects from $51$ categories. We automatically create anomalies for a given 3D shape by applying various deformations to both the geometry and texture. This includes generating realistic cracks and fractures using [39], applying random geometric deformations [43] like bumps, bends, and twists, as well as randomly translating, rotating, and swapping materials in different parts of the shapes. In total, we generated $2345$ anomaly shapes. To generate each scene, we first randomly choose a set of objects consisting of both normal and their anomaly of the same instances. We ensure that the majority of objects in each scene are normal. We also have few scenes where no anomaly is present. The rotational poses for the objects are obtained using rigid body simulation [4]. The objects are scaled and placed into the scene at random locations, ensuring collisions do not occur. We generated total $8K$ scenes. Each scene consists of $3$-$6$ objects rendered in $20$ views. For the training set, we randomly select $5K$ scenes from the $39$ categories. We build two disjoint test sets. The first one (seen) contains $1K$ scenes from the seen categories but with unseen object instances. The second one (unseen) contains $2K$ scenes from rest of the $12$ novel categories.
PartsAD-15K. We use a subset of the ABC dataset [27] that consists of $4200$ shapes. We follow the strategy above to generate anomalies. Additionally, for each shape, we sample geometrically close instances from the dataset and assign them as anomalies to use in the same scene. This approach generates a large set of high-quality anomalies that closely resemble their normal counterparts in high-level geometry, but with subtle shape variations that make them anomalous in the context of the normal ones. In total, we generated $10,203$ anomaly shapes. Using these shapes, we created $15K$ scenes, each consisting of $3$ to $12$ objects rendered from $20$ different viewpoints. We divide the dataset into a $12K$ training set and the rest as the test set. Note that the proposed datasets, source code, and models will be made public based upon publication.
4.2 Implementation details
We use a ResNet50-FPN [30] as our 2D encoder backbone. Our 3D backbone consists of a four-scale encoder-decoder-based 3D CNN [32]. We use $M=5$ images as input, each with a resolution of $256\times 256$. Our architecture is flexible to accept a different $M$ during inference. During training, we consider a total of $2M$ views, which we separate into two sets, each containing $M$ views. We use one set to build the neural volume and the cameras of the other set to render the results, and vice versa. The 3D volume $\bm{F}_{v}$ contains $96\times 96\times 16$ voxels with a voxel size of $4cm$. We sample $128$ points on each ray for rendering. We render the features with a spatial dimension of $32\times 32$. The $I_{t\sigma}$ in Eq. 3 is the ground truth segmentation mask of the input image, we only use it for training. The threshold applied to the density volume $\bm{V}_{\sigma}$ is set to be $0.2$ and we run the DBScan algorithm with its default parameters. We resize the teacher’s (DINOv2) features to the same spatial dimension for loss computation. These features are pre-computed for all scenes using publicly available weights, which are not updated during distillation. These DINOv2 features are then reduced to $d_{f}$=$128$ channel dimensions using PCA before distillation. We employ three sparse voxel attention blocks, and each applies 8-headed attention. The value of $k$ is chosen as $20$. The loss weights $\lambda_{\sigma}$, $\lambda_{f}$, and $\lambda_{p}$ are all set to $1$. We first pretrain the network with only the reconstruction loss for $50$ epochs. Then, we train the network end-to-end with both the reconstruction loss and the binary classification loss for another $50$ epochs. We maintain a batch size of $4$, and use the Adam optimizer with a learning rate $2\times 10^{-5}$. Finally, the run-time of our method is $65$ms on a single A40 GPU for a typical scene with 5 views as input.
4.3 Baseline Comparisons
We quantitatively evaluate the anomaly classification results using two evaluation metrics – the area under the ROC curve (AUC) and accuracy. A prediction is considered correct if the bounding box IoU is greater than 0.5 and the corresponding anomaly classification is correct. We have not designed a separate localization metric as our estimated bounding boxes are very accurate. These metrics are calculated object-wise and then averaged across all test scenes. Since no prior work exists on this task, we define several competitive baselines for comparison. Specifically, in  Tab. 1, we compare our method with two relevant approaches: a reconstruction-based baseline and two multi-view 3D object detection methods. For the reconstruction-based design, we first employ COLMAP [38] to obtain a point cloud reconstruction of the multi-object scene environment. We use default dense reconstruction parameters but utilize the provided ground truth camera matrices. Since 5 views are insufficient, we use a total of 20 views to reconstruct the scene. We extract individual object point clouds from the reconstructed scenes, and train a Siamese style network to obtain their pairwise similarity. We use DGCNN [49] (pretrained on ShapeNet [14]) as the point cloud feature extractor and the triplet loss [22] to supervise the network. Finally, we employ a voting strategy to aggregate the pairwise distances of all objects in the scene and obtain their individual binary labels. We observe that the accuracy of this method is sensitive to the reconstruction quality, leading to poor performance on both datasets. Next, we adapt ImVoxelNet [37] and DETR3D [48], two multi-view object detection frameworks in our problem setting, aiming to locate and classify each object in the scene as either anomaly or normal. ImVoxelNet uses the similar 2D-3D projection as ours to construct the voxel representation and then a 3D detection head [29] outputs the final prediction. DETR3D is a transformer-based design and uses the set prediction loss [9] for end-to-end detection without NMS. We use the ground truth 3D bounding boxes to train these models and the reported scores are calculated object-wise based on the output of the classification head. We observe that while both methods perform well for large cracks or fractures, they struggle when intra-group comparison is necessary. This is because they tend to memorize certain anomaly types without learning to generalize to compare with other objects in the scene. Our method significantly outperforms all baselines (Tab. 1) on both datasets, highlighting the effectiveness of our dedicated architecture for matching corresponding regions. Notably, the performance drop on the unseen set is relatively lower in our case, attributed to the robust 3D representation that effectively generalizes to novel categories. We show qualitative results of our method in Fig. 3
Table 1: Quantitative results. We compare our method to three related works in two datasets and report the results in terms of anomaly detection AUC and accuracy.
Datasets COLMAP [38] ImVoxelNet [37] DETR3D [48] Ours
AUC Accuracy AUC Accuracy AUC Accuracy AUC Accuracy
ToysAD-8K-Seen 73.45 60.48 78.13 65.55 79.16 67.37 91.78 83.21
ToysAD-8K-Unseen 72.86 58.12 73.19 60.12 74.60 62.98 89.15 81.57
PartsAD-15K 72.78 61.34 72.80 64.34 74.49 65.11 86.12 79.68
Figure 3: AD qualitative results on (a) the unseen test categories of ToysAD-8K and (b) the test set of PartsAD-15K using our proposed framework. The green box denotes correct predictions, while the red box indicates where the model misses the anomaly. Due to limited space, one view is shown per scene. Please refer to the supplementary materials for more results.
4.4 Ablations and Model Analysis
We conduct several studies into the performance of the proposed model including changes to architecture, robustness analysis and real-world experiments. Unless stated otherwise, all experiments in this section are carried out on the ToysAD-8K unseen set using 5 input views.
Figure 4: Correspondences are obtained in 3D space using the neural feature field $\bm{V}_{f}$, then projected onto 2D views using camera matrices for visualization. The feature field is rendered at respective viewpoints, the first 3 PCA components are used, with each matched to a different color channel.
Ablation of architecture design. We ablate the core components in our model and report the results in Tab. 2. All variants include the 3D feature fusion module and are optimized at least for the image reconstruction loss, which is essential for constructing scene geometry. Variant A directly maps object-centric features to their binary labels using an MLP without comparing them, while for variant B, standard attention layers are applied to the object-centric features to learn cross-object correlations. Despite the attention layers, B only shows a slight improvement (+2.1% AUC) over A.
Table 2: Ablation Results on ToysAD-8K
Variants AUC Accuracy
A: baseline 79.13 66.78
B: A + vanilla attention 81.24 68.20
C: B + feature distillation 87.05 79.56
Final: C + sparse voxel attention 89.15 81.57
We then introduce DINOv2 feature distillation (variant C), which significantly boosts performance by 5.8% AUC and 11.3% accuracy, indicating the importance of the part-aware 3D representation and correspondences for our task (also see Fig. 4). Our final design achieves the performance gain by utilizing sparse voxel attention, which focuses on the top-$k$ most relevant features. This leverages robust correspondences learned through DINOv2 feature distillation, effectively eliminating noisy correlations and directing attention solely to corresponding object regions.
Robustness. Our method achieves some robustness to occlusion by effectively using input from multiple viewpoints. Fig. 7 illustrates examples where our model accurately identifies the anomalous region, even when occluded in some views. In Fig. 5 (left), we show how well our model can perform with additional or reduced input views at test time. We train our model with 5 views and then evaluate it using 1, 3, 5, 10 and 20 views. These results clearly indicate that our model can perform reasonably well with only 3-5 views; however, additional views can be used to boost performance at test time.
Figure 5: Impact of number of views (left) and object count (right) on model performance.
We analyze the impact of object count on model performance (see Fig. 5 (right)) using the PartsAD-15K dataset. Increasing the number of objects in a scene usually leads to more occlusion and lower resolution for each individual object. However, the performance drop between the two extremes is minimal ($<$ 1.5% AUC) as shown in the figure. In addition, we investigate the model’s ability to generalize to more objects at test time than it observed during training. To this end, we trained our model on scenes with 3-7 objects and tested on two sets: one with 3-7 objects (AUC: 86.75) and another with 8-12 objects (AUC: 85.10). These results demonstrate our model’s adaptability to varying object counts.
In another study, we test our primary objective of choosing the majority group as normal and the rest as anomalies using an example shown in Fig. 6. We create five scenes (a-e) by using two geometrically similar objects, and gradually introduce more from one (1 to 5 respectively) while maintaining a total of six objects. For example, in scene (a), the first object appears only once, making it an anomaly. Similarly, in scene (e), the second object appears only once, hence considered an anomaly. We utilize a consistent background across all scenes to ensure uniformity. As shown in the figure, our model correctly classifies the anomalies in each scene. We note scene (c) presents an ambiguous case, where both objects appear in equal numbers. Despite this, our model is able to separate the two groups.
Real world testing. Here we apply our model, which is trained on the synthetic dataset, on a small set of real test scenes. Each scene is set up in an indoor environment with adequate lighting and is captured using a 3D scanning software [1]. This results in a set of input views with globally optimized cameras. Fig. 8 illustrates the results for three such scenes.
Limitations. Our benchmark and model also have a few limitations. Firstly, this work focuses solely on a limited set of anomalies that are common to manufacturing scenarios; potentially missing other anomalies in real-world scenarios. Since acquiring real damaged objects is expensive and difficult, our dataset primarily uses shapes of synthetic objects. Moreover, our model assumes that object instances are rigid and cannot handle articulations or deformations. It also assumes that objects are not touching or completely occluded in space. Moreover, a scene with mostly anomalies can be challenging for our model due to the lack of ‘normal’ data points for comparison (except for fractures or cracks). Moreover, the performance is dependent on the anomalous region being captured in at least one view. Finally, in real-world scene testing, the noisy camera poses as well as the gap between synthetic and real-world environments may degrade the performance.
Figure 6: We create five scenes (a-e) with two similar objects, varying their counts per scene while maintaining a constant total. Our model correctly identifies minorities as anomalies in all cases except (c), where the equal number of objects creates ambiguity. Our model still selects one group.
Figure 7: Resolving (a) occlusion and (b) 3D ambiguity using multi-view images. The anomaly ‘sheep’ in (a) has a missing tail (only visible in the last view due to occlusion), and the ‘hammer’ handle in (b) is bent (only apparent from the last view-angle due to 3D ambiguities).
Figure 8: Three real-world scenes are tested using our method and it can successfully detect all anomalous instances. Five input views are used for testing, while one view for each scene is used for visualization.
5 Conclusion
In this paper, we have introduced a novel AD problem inspired by real-world applications along with two new benchmarks. The proposed task goes beyond the traditional AD setting and involves a cross study of objects in a scene from multiple camera viewpoints to identify the ‘odd-looking’ minority group. We show that our model is robust to varying number of views and objects, and outperforms the baselines that do not consider cross-object correlations.
Broader Impacts.
The proposed techniques could be potentially used in improving quality checks and product safety by providing warnings to the users in production lines. The authors are not aware of any potential harm that may arise when the technology is used.
Appendix A Appendix / supplemental material
This appendix is structured as follows: we present additional details of the proposed datasets in Sec. A.1, training details in Sec. A.2, and additional qualitative results in Sec. A.3.
A.1 Data Generation Details
Our proposed scene AD datasets, ToysAD-8K and PartsAD-15K, are built upon two publicly available 3D shape datasets: Toys4K [41] (Creative Commons and royalty-free licenses) and ABC [27] (MIT license). For the ToysAD-8K, we selected 1,050 shapes from the Toys4K dataset, focusing on the most common real-world objects across 51 categories. A complete list of these categories is provided in Table 3. On the other hand, PartsAD-15K is a non-categorical dataset. For this dataset, we randomly selected a subset of 4,200 shapes from the large-scale ABC.
For both datasets, we consider the following anomaly types: cracks, fractures, geometric deformations (e.g., bumps, bends, and twists), translation, rotation, material mismatch, and missing parts. Additionally, for PartsAD-15K, we use a different but geometrically similar instance in the same scene as an anomaly. To obtain a geometrically similar shape, we use feature-based KNN clustering. Specifically, we extract DINOv2 features from multi-view images (rendered from multiple fixed viewpoints) of the 3D shapes, concatenate these features, and use the resulting concatenated features to build the KNN cluster. Then, we retrieve a similar 3D shape by querying the KNN cluster. To ensure the retrieved shape is geometrically similar, we calculate the Chamfer distance between the shapes and only accept a shape if the distance is below a certain threshold.
Table 3: Dataset composition of ToysAD-8K
Categories
Seen dinosaur, fish, frog, monkey, light, lizard, orange, boat, dog, lion, pig, cookie, panda, chicken,
orange, ice, horse, car, airplane, cake, shark, donut, hat, cow, apple, bowl, hamburger, octopus,
giraffe, chess, bread, butterfly, cupcake, bunny, elephant, fox, deer, bus, bottle
Unseen mug, plate, robot, glass, sheep, shoe, train, banana, cup, key, penguin, hammer
Our anomaly generation process is automatic. To ensure the quality of the generated anomalies, we perform several checks. For example, during positional or rotational anomaly creation, if a part detaches from the main body during deformation, we reject the anomaly and try again with adjusted parameters. Similarly, if removing a part makes the shape impractical, we discard it and try removing a different part. For fracture anomaly, if a particular fracture removes more than 90% or less than 10% of an object we discard the sample and regenerate another fracture. Finally, we ensure the anomalous region of an object is visible from at least one viewpoint.
To generate a realistic scene environment, we use PBR materials [44] for floors and HDRI environment maps [56] for image-based lighting to illuminate scenes. We randomly select a pair of PBR material and HDRI environment maps from the assets to randomize the scene background. Objects are placed randomly such that no collisions occur, and each object’s rotational poses are obtained using Blender’s rigid body simulation [43]. We employed Blender 2.93 [43] with Cycles ray-tracing renderer for photo-realistic rendering. Blender 2.93 is released under the GNU General Public License (GPL, or “free software”), and the PBR and HDRI maps are released under the CC0 license.
Our framework can easily be trained with real-world manufacturing scene environments. Our model relies solely on 2D supervision, making the data collection process much easier. We also do not need precise annotation of 3D bounding boxes as they are not used during training. For annotating instance-wise anomaly labels, we can use 2D bounding boxes, which can be projected in 3D using Visual Hull [28], then used for coarse localization.
A.2 Training Details
We train our model in two stages. In the first stage, we train it with just image and feature reconstruction losses. In the second stage, we train the model end-to-end with both reconstruction and binary classification losses. All experiments are performed on a single NVIDIA A40 GPU with a batch size of 4, utilizing 28GB of GPU memory. The first stage takes 36 hours to complete, followed by an additional 24 hours for the second stage.
The ablation experiments are conducted on the same workstation with the same GPU by removing one or a few core components from the full method. Specifically, variant methods A and B take $36$ hours to train the first stage, while only taking $14$ hours to train the second stage. Regarding variant method C, it takes similar $36$ and $24$ hours for the two stages as in the full method.
We compare our method with COLMAP (BSD license), ImVoxelNet (MIT license), and DETR3D (MIT license). For the COLMAP-based approach, we use DGCNN (MIT license) as a feature extractor. In terms of training time of these baseline methods, COLMAP takes $10$ hours to train, ImVoxelNet takes $24$ hours to train, and DETR3D takes $2$ days to converge.
The standard deviation of all our experiments (including the ablations and our method) under multiple runs is less than 0.5.
A.3 Additional Qualitative Results
In Fig. 9 and Fig. 10, we present additional qualitative results on ToysAD-8K and PartsAD-15K, respectively.
Figure 9: Additional results on the unseen set of ToysAD-8K dataset. For rows 1 to 3, the anomalies are easy to spot and self-explanatory. In row 4, there are two anomalies: one with broken outer parts at the back (see view 5), and another with a tilted roof (see views 1 and 2). For row 5, one leg is broken (see view 2). In row 6, the eye is missing (see view 5).
Figure 10: Additional results on the PartsAD-15K dataset.

### v2

@@
 Abstract
-This paper introduces a novel anomaly detection (AD) problem that focuses on identifying ‘odd-looking’ objects relative to the other instances within a scene.
-Unlike the traditional AD benchmarks, in our setting, anomalies in this context are scene-specific, defined by the regular instances that make up the majority.
-Since object instances are often partly visible from a single viewpoint, our setting provides multiple views of each scene as input.
+This paper introduces a novel anomaly detection (AD) problem that focuses on identifying ‘odd-looking’ objects relative to the other instances in a given scene.
+In contrast to the traditional AD benchmarks, anomalies in our task are scene-specific, defined by the regular instances that make up the majority.
+Since object instances may be only partly visible from a single viewpoint, our setting employs multiple views of each scene as input.
 To provide a testbed for future research in this task, we introduce two benchmarks, ToysAD-8K and PartsAD-15K.
-We propose a novel method that generates 3D object-centric representations for each instance and detects the anomalous ones through a cross-examination between the instances.
-We rigorously analyze our method quantitatively and qualitatively in the presented benchmarks.
+We propose a novel method that constructs 3D object-centric representations from multiple 2D views for each instance and detects the anomalous ones through a cross-instance comparison.
+We rigorously analyze our method quantitatively and qualitatively on the presented benchmarks.
+The datasets, source code, and models will be made available upon publication.
+Figure 1: (a) We propose a new anomaly detection task focused on identifying ‘odd-looking’ objects relative to other instances within a scene.
+Inspired by real-world quality control in production environments, this task aims to detect subtle variations in geometry and texture, including defects like cracks and fractures, in a group of manufactured samples.
+(b) Our setting is scene-specific, requiring a comparison of object instances within the input scene, unlike the standard AD setting, which takes only a single object as input.
+(c) Our matching-based paradigm enables cross-category performance.
 1 Introduction
 Anomaly detection (AD) [10, 29] aims to detect patterns that deviate from expected behavior.
-In standard computer vision AD benchmarks, the non-conforming (or anomalous) patterns in images are often either due to high-level variations such as introduction of an object instance from an unseen category [2, 8, 12] or low-level variations in object shape and texture [6, 16] (see Fig.
+In standard AD benchmark for computer vision, anomalies may arise from high-level factors such as introduction of an object instance from an unseen category [9, 31, 2] or low-level ones such as defects and variations in the normal object shape and texture [8, 6, 13].
+In these benchmarks, AD methods analyze each object individually, ignoring relationships between objects in the same scene or contextual factors.
+However, in many real-world scenarios, ‘normality’ can be conditional.
+For instance, in a production line for blue-handled coffee cups, a yellow handle is anomalous, and in a production line for yellow-handled coffee cups, a blue handle is anomalous (see Fig.
+1(b)).
+This context dependent normality cannot be addressed by methods assuming a fixed standard of normality.
+Inspired by real-world visual inspection, we introduce a new AD problem with two new benchmarks and a novel solution.
+In our setup (see Fig.
+1(a)), each image contains a scene with multiple instances of the same object (e.g., coffee cup) including either all normal, or a mix of normal and anomalous instances.
+Since defining exact shape and appearance specifications on paper is often impractical, we assume each scene has multiple normal instances together providing a scene-specific reference for ‘normality’.
+Our objective is to detect anomalous instances in a scene and to generalize to previously unseen scenes including unseen objects and layouts.
+Unlike existing benchmarks, this task often requires a holistic understanding of the scene by comparing instances with each other with the exception of few anomalies such as cracks and misaligned parts that can be identified by inspecting individual instances.
+Additionally, since performing AD from a single view can be ambiguous due to self-occlusion and occlusion between the instances, we provide multiple views of the scene to cover its entire relevant extent unlike existing benchmarks with single-view inputs (see Fig.
 1(a)).
-In these benchmarks, the definitions of normal and anomalous patterns are typically not predefined but implicitly learned through representations and/or classifiers [21, 2, 16] that are invariant to appearance changes within the normal data distribution while being sensitive to the anomalous ones.
-However, in many real-world applications such as quality control in production lines, product specification, hence the definition of ‘normality’, is not only known in advance but also specific to each object instance.
-For example, while a coffee cup with a red handle is considered normal and one with a blue handle is anomalous when the goal is to produce red-handled cups, but the opposite is true if the goal is to produce blue-handled cups.
-This instance-specific definition of normality cannot be addressed by methods that rely on a global and implicit definition of normality.
-In this paper, inspired from real-world visual inspection scenarios, we introduce a new AD problem, along with two new benchmarks with different characteristics and a novel solution.
-As depicted in Fig.
-1 (b), in this proposed setting, each image contains a scene with multiple instances of the same object (e.g., coffee cups) including either only normal, or a mix of normal and anomalous instances where the anomalous instances constitute the minority.
-As defining precise visual specifications for shape and appearance is often infeasible, we assume that normal instances, which form the majority, provide a scene-specific reference for ‘normality’.
-Our objective is to detect anomalous instances in a scene while generalizing to previously unseen scenes including novel objects and spatial configurations.
-Unlike existing benchmarks, our task often requires a holistic understanding of the scene by comparing instances with each other, although some anomalies, such as cracks and misaligned parts, can be identified by inspecting individual instances.
-Additionally, since performing AD from a single view can be ambiguous due to self-occlusion and occlusion between the instances, we provide multiple views of the scene to cover its entire relevant extent, unlike the existing benchmarks that provide a single view as input (see Fig.
-1(c)).
-Our goal is to detect anomalous samples in previously unseen scenes from multiple views.
-Figure 1: The standard and our new AD settings.
-The proposed problem presents several challenges requiring the following capabilities: i) 3D understanding of the scene and registering the views from multiple camera viewpoints without groundtruth 3D knowledge while identifying potential occlusions, ii) aligning and comparing object instances with each other without their relative pose information in both training and evaluation, iii) learning representations that generalize to unseen object instances during testing.
-To address these challenges, we propose a novel method that takes as input multiple views of the same scene, projects them into a 3D voxel grid, produces a 3D object-centric representation for each instance, and predicts their labels by cross-correlating instances through an efficient attention mechanism.
-We leverage recent advances in differentiable rendering [31] and self-supervised learning [33] to supervise 3D representation learning.
+The proposed problem presents several challenges requiring the following key capabilities: i) achieving 3D understanding of the scene and view registration from multiple camera viewpoints without groundtruth 3D knowledge while handling occlusions, ii) aligning and comparing object instances without their pose information during training and testing, iii) learning representations that generalize to unseen object instances from seen or unseen categories at test time.
+To tackle these challenges, we introduce a novel method that takes as input multiple views of the same scene, projects them into a 3D voxel grid, generates a 3D object-centric representation for each instance, and predicts their labels by cross-correlating instances using an efficient attention mechanism.
+We utilize advances in differentiable rendering [26] and self-supervised learning [28] to supervise 3D representation learning.
 Specifically, we render the voxel representations for multiple viewpoints and match them with the given 2D views to produce geometrically consistent features.
-Additionally, we encourage our model to learn part-aware 3D representations by distilling features from 2D self-supervised model DINOv2 [33], enhancing the correspondence matching across instances.
-Finally, since there is no prior benchmark for this task, we propose two new benchmarks – ToysAD-8K and PartsAD-15K, including instances from common semantic object categories as well as mechanical parts respectively to provide a testbed for future research.
-Our method significantly outperforms various baselines that do not compare instances with each other, and we rigorously analyze various aspects of our model and benchmarks.
+Additionally, we encourage our model to learn part-aware 3D representations by distilling features from 2D self-supervised model DINOv2 [28], enhancing the correspondence matching across instances and their alignment.
+Finally, since there is no prior benchmark for this task, we propose two new benchmarks – ToysAD-8K and PartsAD-15K, which include common objects categories as well as mechanical parts respectively to facilitate further research.
+Our method significantly outperforms baselines that do not compare instances with each other, and we rigorously analyze various aspects of our model and benchmarks.
 2 Related Work
@@
 Few-shot AD.
-There are some works [18, 23, 52, 50, 7] aiming to detect anomalies from a small number of normal samples as support images.
+There are some works [15, 19, 46, 44] aiming to detect anomalies from a small number of normal samples as support images.
 In our setting, the concept of normality in each image is also learned from a few instances only.
@@
 Similarly, in this work, we focus on learning a feedforward model in a multi-object scene environment using sparse multi-view images but for AD.
-Leveraging foundation models.
+Figure 2: Overview of our framework.
+We extract features from a sequence of input views using a 2D CNN and back-project them into a 3D volume, which is then refined with a 3D CNN, resulting in $\bm{F}_{v}$.
+Next, we extract object-centric feature volumes $\{\bm{z}_{n}\}_{n=1}^{N}$, which are fed into the cross-instance matching module to learn correlations among objects using sparse voxel attention.
+To improve the 3D representation of the scene, we distill the knowledge of a 2D vision model namely DINOv2, and integrate the learned knowledge into our 3D network via differentiable rendering.
+This aids in obtaining a part-aware and geometrically consistent 3D feature representation.
+Leveraging large-scale models.
 Large-scale pretraining on image datasets has shown impressive generalization capabilities on various tasks [30, 21, 28].
-Previous works [57, 3] demonstrate that features extracted from DINOv2 [33] serve as effective dense visual descriptors with localized semantic information for dense correspondence estimation task.
-Some recent works [26, 46] use feature distillation techniques to leverage 2D foundation vision models for 3D task.
-Inspired by these works we utilize DINOv2 to distill its dense semantic knowledge into our 3D network that enables our network to infer robust local correspondences, which aids in fine-grained object matching.
+Previous works [50, 3] demonstrate that features extracted from DINOv2 [28] serve as effective dense visual descriptors with localized semantic information for dense correspondence estimation tasks.
+Some recent works [22, 40] use feature distillation techniques to leverage 2D foundation vision models for 3D tasks.
+Inspired by these works, we utilize DINOv2 to distill its dense semantic knowledge into our 3D network enabling our network to infer robust local correspondences, which aids in fine-grained object matching.
 3 Method
-3.1 Overview
-Consider a scene containing multiple rigid objects $\{o_{n}\}_{n=1}^{N}$ of the same instance, where $N$ represents the number of objects in the scene.
-Each object’s pose is arbitrary and unknown.
+Consider a scene containing $N$ instances $\{o_{n}\}_{n=1}^{N}$ of the same rigid object where their poses are unknown.
 The goal is to identify anomalies in the group of objects by assessing their mutual similarity.
-An observation in our setting consists of $M$-view $H\times W$ dimensional RGB images =$\mathcal{I}=\{\bm{I}_{t}\}_{t=1}^{M}$, and their corresponding camera projection matrices $\mathcal{P}=\{\bm{P}_{t}\}_{t=1}^{M}$, $\bm{P}_{t}=\bm{K}\left[\bm{R}_{t}|\bm{T}_{t}\right]$, with intrinsic $\bm{K_{t}}$, rotation $\bm{R}_{t}$ and translation $\bm{T}_{t}$ matrices.
-In our setting, $M$ is $5$, forming sparse-view inputs.
+An observation in our setting consists of $M$-view $H\times W$ dimensional RGB images =$\mathcal{I}=\{\bm{I}_{t}\}_{t=1}^{M}$, and their corresponding camera projection matrices $\mathcal{P}=\{\bm{P}_{t}\}_{t=1}^{M}$ where $\bm{P}_{t}=\bm{K}\left[\bm{R}_{t}|\bm{T}_{t}\right]$, with intrinsic $\bm{K_{t}}$, rotation $\bm{R}_{t}$ and translation $\bm{T}_{t}$ matrices.
+We set $M$ to $5$, forming sparse-view inputs in our experiments.
 Our goal is to learn a mapping $\psi$ from the multi-view images to object-centric anomaly labels ${y}_{n}\in\{0,1\}$ and its corresponding 3D bounding box $\bm{b}_{n}$, defined as:
 $\psi:\{(\bm{I}_{t},\bm{P}_{t})\}_{t=1}^{M}\mapsto\{({y}_{n},\bm{b}_{n})\}_{n=1}^{N}.$ (1)
-Note, labels $y_{n}$ are defined relative to other objects in the scene.
-For example, consider a group of three coffee cups, of which two have red handles, and one has a blue handle.
+Note that labels $y_{n}$ are defined relative to other objects in the scene.
+For example, consider a group of three coffee cups, of which two have yellow handles, and one has a blue handle.
 In this example, the latter cup is considered an anomaly.
-Figure 2: Overview of our framework.
-We extract features from a sequence of input views using a 2D CNN and then back-project them into a 3D volume, which is then refined using a 3D CNN, resulting in $\bm{F}_{v}$.
-We then extract object-centric feature volumes $\{\bm{z}_{n}\}_{n=1}^{N}$ using RoI pooling, where $\mathcal{B}$ estimates the RoI region of each object.
-The cross-instance matching module $\mathcal{M}$ then learns a correlation among the objects using a sparse voxel attention.
-To improve the 3D representation of the scene, we distill the knowledge of a 2D vision model namely DINOv2, and integrate the learned knowledge into our 3D network via differentiable rendering.
-Our architecture, illustrated in Fig.
-2, is comprised of three main components: First, the 3D feature fusion module encodes each view image and projects it to 3D, forming a fused 3D feature volume.
-Second, the feature distillation block is employed to enhance the 3D feature volume through differentiable rendering.
-This facilitates finding local correspondences between objects.
-Finally, the cross-instance matching module leverages established correspondences to compare all similar object regions in the scene using a sparse voxel attention mechanism.
+Our pipeline, illustrated in Fig.
+2, is comprised of four main components: First, the 3D feature volume construction module (Sec.
+3.1) encodes each view image and projects it to 3D, forming a fused 3D feature volume.
+Second, the feature enhancement block (Sec.
+3.2), employed only during training, facilitates the learning of a part-aware and geometrically consistent 3D space through differentiable rendering and feature distillation.
+Third, the object-centric feature extraction module (Sec.
+3.3) locates each object in the scene and represents it with a feature volume.
+Finally, the cross-instance matching module (Sec.
+3.4) efficiently compares all similar object regions using a sparse voxel attention mechanism and predicts instance anomalous labels and 3D coordinates.
 Next, we elaborate on details.
 3.1 3D Feature Volume Construction
-We first extract 2D features $\bm{F}_{t}=\mathcal{E}_{2D}(\bm{I}_{t})\in\mathbb{R}^{d\times h\times w}$ for each input view using a shared CNN encoder $\mathcal{E}_{2D}$, where $d$ is the feature dimension.
+In our first module, we obtain a 3D feature volume representation given multi-view images as input.
+Specifically, we first extract 2D features $\bm{F}_{t}=\mathcal{E}_{2D}(\bm{I}_{t})\in\mathbb{R}^{d\times h\times w}$ for each input view using a shared CNN encoder $\mathcal{E}_{2D}$, where $d$ is the feature dimension.
 These 2D features are then projected into 3D voxel space as follows:
 $\bm{F}_{v}=\mathcal{E}_{3D}(\texttt{aggr}(\{\Pi_{\texttt{proj}}(\bm{F}_{t},\bm{P}_{t})\}_{t=1}^{M})),$ (2)
-where $\Pi_{\texttt{proj}}$ back-projects each view feature $\bm{F}_{t}$ using known camera intrinsic and extrinsic, generating 3D feature volumes of size $d\times v_{x}\times v_{y}\times v_{z}$.
+where $\Pi_{\texttt{proj}}$ back-projects each view feature $\bm{F}_{t}$ using known camera intrinsic and extrinsic parameters, generating 3D feature volumes of size $d\times v_{x}\times v_{y}\times v_{z}$.
 These feature volumes are aggregated over all input views using an average operation as in [27, 37].
-Finally, a 3D CNN-based network $\mathcal{E}_{3D}$ is employed to refine the aggregated feature volume, resulting in a final voxel representation $\bm{F}_{v}$.
+Finally, a 3D CNN-based network $\mathcal{E}_{3D}$ is employed to refine the aggregated feature volume, resulting in a final voxel representation $\bm{F}_{v}$ of size $d\times v_{x}\times v_{y}\times v_{z}$, which is further fed into the feature enhancement block, described next, to obtain a part-aware and geometrically consistent 3D representation.
+3.2 Feature Enhancement Block
 We use volume rendering [26] to reconstruct the geometry and appearance of the scene.
@@
 Formally, the image rendering and its corresponding loss function for a single viewpoint $\bm{P}_{t}$ are shown below:
-$[\hat{\bm{I}}_{t},\hat{\bm{I}}_{t\sigma}]=\mathcal{R}([\bm{V}_{c},\bm{V}_{\sigma}],\bm{P}_{t}),\quad\mathcal{L}^{\text{im}}_{t}=||\bm{I}_{t}-\hat{\bm{I}}_{t}||^{2}+\lambda_{\sigma}||\bm{I}_{t\sigma}-\hat{\bm{I}}_{t\sigma}||^{2},$ (3)
-where $\hat{\bm{I}}_{t}$ and $\hat{\bm{I}}_{t\sigma}$ are rendered image and mask. $\lambda_{\sigma}$ is a loss weight.
-We propose to improve the 3D voxel representation $\bm{F}_{v}$ by also reconstructing neural features instead of just color and density.
+$\begin{split}\mathcal{L}^{\text{im}}_{t}=||\bm{I}_{t}-\hat{\bm{I}}_{t}||^{2}+||\bm{I}_{t\sigma}-\hat{\bm{I}}_{t\sigma}||^{2},\end{split}$ (3)
+where $[\hat{\bm{I}}_{t},\hat{\bm{I}}_{t\sigma}]=\mathcal{R}([\bm{V}_{c},\bm{V}_{\sigma}],\bm{P}_{t})$, $\hat{\bm{I}}_{t}$ and $\hat{\bm{I}}_{t\sigma}$ are the rendered image and mask.
+Moreover, we improve the 3D voxel representation $\bm{F}_{v}$ by reconstructing neural features in addition to geometry and appearance.
 We supervise the feature reconstruction by a pretrained 2D image encoder $\Phi$ as a teacher network.
 We choose DINOv2 [28] as the teacher network due to its excellent ability to capture various object geometries and correspondences.
-We use a projector function $\beta$, implemented as a four-layered $1\times 1\times 1$ convolution block that projects $\bm{F}_{v}$ to a neural feature field $\bm{V}_{f}$, changing the channel dimension from $d$ to $d_{f}$.
-Similar to color rendering, we generate rendered features $\hat{\Phi}_{t}$ of size $d_{f}\times h_{f}\times w_{f}$ at a given viewpoint using volume renderer $\mathcal{R}$.
-The objective is to minimize the difference between the rendered features and the teacher’s features $\Phi(\bm{I}_{t})$.
-We choose cosine distance as our feature loss ($\mathcal{L}^{\text{feat}}_{t}$) which we find easier to optimize compared to the standard L2 loss.
-We apply stop-gradient to density in rendering of features $\hat{\Phi}_{t}$, as the teacher’s features are not fully multi-view consistent [3], which could harm the quality of reconstructed geometry.
-The final reconstruction loss is the sum of all image and feature reconstruction losses weighed by a factor $\lambda_{f}$:
-$\mathcal{L}^{r}=\sum_{t=1}^{M}(\mathcal{L}^{\text{im}}_{t}+\lambda_{f}\mathcal{L}^{\text{feat}}_{t}).$ (4)
+To this end, we use a projector function $\beta$, instantiated as a four-layered $1\times 1\times 1$ convolution block that projects $\bm{F}_{v}$ to a neural feature field $\bm{V}_{f}$, changing the channel dimension from $d$ to $d_{f}$.
+Similar to color rendering, taking $\bm{V}_{f}$ and $\bm{V}_{\sigma}$ as input, we generate rendered features $\hat{\Phi}_{t}$ of size $d_{f}\times h_{f}\times w_{f}$ at a given viewpoint using volume renderer $\mathcal{R}$, and minimize the difference between the rendered features and the teacher’s features $\Phi(\bm{I}_{t})$.
+We choose cosine distance as our feature loss ($\mathcal{L}^{\text{feat}}_{t}$), which we find easier to optimize compared to the standard L2 loss.
+We do not allow the gradients to follow through the density in the rendering of features $\hat{\Phi}_{t}$, as the teacher’s features are not fully multi-view consistent [3], which could harm the quality of reconstructed geometry.
+The final reconstruction loss is the sum of all image and feature reconstruction losses:
+$\mathcal{L}^{r}=\sum_{t=1}^{M}(\mathcal{L}^{\text{im}}_{t}+\mathcal{L}^{\text{feat}}_{t}).$ (4)
 The key benefits of reconstructing DINOv2 features are twofold.
-1) Firstly, distilling features from general-purpose feature extractors pre-trained on large external datasets, incorporates open-world knowledge into the 3D representation.
+First, distilling features from general-purpose feature extractors pretrained on large external datasets incorporates open-world knowledge into the 3D representation.
 This enables our model to perform significantly better on unseen object instances or even on novel categories, as demonstrated in the experiments.
-2) Secondly, the distillation enforces consistent 3D scene representation, leading to identical features for the same object geometries.
+Second, the distillation through rendering enforces consistent 3D scene representation, leading to very similar features for the same object geometries.
 This enables the model to infer robust local correspondences (see Fig.
-4), which aids in fine-grained object matching.
-3.3 Object-centric 3D Features Extraction
-We extract bounding box regions of objects using a box estimator $\mathcal{B}$.
-First, we obtain a voxel reconstruction of the scene by applying a threshold to the predicted density $\bm{V}_{\sigma}$.
-Then, we employ DBScan [20], a density-based clustering method to retrieve all bounding box regions $\{\bm{b}_{n}\}_{n=1}^{N}$ corresponding to the objects in the scene.
-Next we obtain object-centric feature volumes $\{\bm{z}_{n}\}_{n=1}^{N}$, each with a size of $8\times 8\times 8$ from these regions through applying RoI pooling [17].
-To extract correspondence indices given two objects with feature volumes $\bm{z}_{n}$ and $\bm{z}_{m}$, We formally define a function $\mathcal{C}_{k}$:
-$\mathcal{C}_{k}(\bm{z}_{n},\bm{z}_{m})=\texttt{top}_{k}[\beta(\bm{z}_{n})^{T}\beta(\bm{z}_{m})].$ (5)
-The function returns top-$k$ most relevant feature locations in $\bm{z}_{m}$ for each voxel location in $\bm{z}_{n}$.
-This is achieved by first projecting each voxel feature using the same projector function $\beta$ and then computing their voxel-level pairwise similarity.
-We use $\mathcal{C}_{k}$ to perform sparse attention-based comparisons between multiple object volumes, as described next.
+6), which aids in fine-grained object matching.
+3.3 Object-centric Feature Extraction
+Having the enhanced 3D feature volume representation, we aim to localize each object instance in the scene and represent each with a 3D feature volume such that the instances can be compared to each other in the next step.
+To this end, we use the predicted density volume $\bm{V}_{\sigma}$, obtained in Sec.
+3.2, to extract a coarse point-cloud structure of the scene by applying a threshold to $\bm{V}_{\sigma}$.
+Then, we employ DBScan [17], a density-based clustering method to segment the foreground objects in the point cloud.
+This produces bounding box regions $\{\bm{b}_{n}\}_{n=1}^{N}$ corresponding to the objects in the scene.
+Finally, we apply RoI pooling [14] on the voxel representation $\bm{F}_{v}$ to extract object-centric feature volumes $\{\bm{z}_{n}\}_{n=1}^{N}$, each with a size of $d\times 8\times 8\times 8$.
+Note that we do not require ground truth box coordinates during training.
+Next, we pass them to the cross-instance matching module for fine-grained comparison.
 3.4 Cross-instance Matching
-The cross-instance matching module $\mathcal{M}$ takes the object features $\{\bm{z}_{n}\}_{n=1}^{N}$, effectively learns to correlate them using sparse voxel attention, and subsequently predicts their object-specific labels.
+Our task requires comparing object instances with each other, which further demands geometric alignment for each pair.
+Since the relative pose of each instance with respect to the camera is unknown, we avoid explicit alignment and instead focus on finding local matches between them.
+To this end, for a given pair of object instances ($m$ and $n$), we compute the most similar local features between their feature volumes ($\bm{z}_{n}$ and $\bm{z}_{m}$) respectively:
+$\mathcal{C}^{nm}_{k}[i]=\texttt{top}_{k}[\beta(\bm{z}_{n}[i])^{T}\beta(\bm{z}_{m}[.])],$ (5)
+where we denote the operation as $\mathcal{C}^{nm}_{k}$, which returns top-$k$ most relevant feature locations in $\bm{z}_{m}$ for the $i$-th voxel location in $\bm{z}_{n}$.
+This is achieved by first projecting each voxel feature utilizing the projector function $\beta$ (described in Sec.
+3.2 and then computing their voxel-level pairwise similarity followed by a $\texttt{top}_{k}$ operation.
+We use $\mathcal{C}^{nm}_{k}$ to perform sparse attention-based comparisons between multiple object volumes, as described next.
+The query, key, and value embeddings are calculated using linear projections as:
+$\begin{split}\bm{Q}_{n}[i]=\bm{W}^{Q}{\bm{z}}_{n}[i],\quad\bm{K}_{n}[i]=\bm{W}^{K}{\bm{z}}_{n}[i],\\ \quad\bm{V}_{n}[i]=\bm{W}^{V}{\bm{z}}_{n}[i],\end{split}$ (6)
+where the weights $\bm{W}^{Q}$, $\bm{W}^{K}$ and $\bm{W}^{V}$ are shared across all objects. Then, the attention is calculated as:
+$\bar{\bm{z}}_{n}[i]=\sum_{\begin{subarray}{c}m=1\\ m\neq n\end{subarray}}^{N}\sum_{j\in\mathcal{C}_{k}^{nm}[i]}\text{softmax}\left(\frac{\bm{Q}_{n}[i]\bm{K}_{m}[j]}{\sqrt{d}}\right)\bm{V}_{m}[j].$ (7)
 Unlike the vanilla self-attention module in standard transformers [16] that uses all tokens for the attention computation, which is inefficient for our task and may introduce noisy interactions with irrelevant features, potentially degrading performance.
-To overcome this, we compute the sparse voxel attention only among geometrically corresponding voxel locations, as shown in the inset.
-Let $z_{n}[i]\in\mathbb{R}^{d}$ denote the $i$-th voxel of the $n$-th object volume in the scene. The query, key, and value embeddings are calculated using linear projections as:
-$\bm{Q}_{n}[i]=\bm{W}^{Q}{\bm{z}}_{n}[i],\quad\bm{K}_{n}[i]=\bm{W}^{K}{\bm{z}}_{n}[i],\quad\bm{V}_{n}[i]=\bm{W}^{V}{\bm{z}}_{n}[i],$ (6)
-where the weights $\bm{W}^{Q}$, $\bm{W}^{K}$ and $\bm{W}^{V}$ are shared across all objects.
-Let $c_{k}^{nm}[i]$ denote the set of all corresponding voxel indices obtained using Eq. 5. Then, the attention is calculated as:
-$\bar{\bm{z}}_{n}[i]=\sum_{\begin{subarray}{c}m=1\\ m\neq n\end{subarray}}^{N}\sum_{j\in c_{k}^{nm}[i]}\text{softmax}\left(\frac{\bm{Q}_{n}[i]\bm{K}_{m}[j]}{\sqrt{d}}\right)\bm{V}_{m}[j].$ (7)
-The updated feature volume $\bar{\bm{z}}_{n}$ is passed through 3D CNN blocks to downsample by a factor of $1/8$, which is finally reshaped into a vector and fed to a 2-layer MLP outputting the final prediction $\hat{y}_{n}$.
+To overcome this, we compute the sparse voxel attention only among geometrically corresponding voxel locations using $\mathcal{C}_{k}^{nm}$.
+Next, the updated feature volume $\bar{\bm{z}}_{n}$ is passed through 3D CNN blocks to downsample by a factor of $1/8$, which is finally reshaped into a vector and fed to a 2-layer MLP outputting the final prediction $\hat{y}_{n}$.
 The classification loss is calculated as:
 $\mathcal{L}^{\text{bce}}=\sum_{n=1}^{N}\ell_{\text{bce}}(\hat{y}_{n},y_{n}),$ (8)
-where $\ell_{\text{bce}}$ is the binary cross-entropy loss function.
-The total training loss of our framework is $\mathcal{L}=\mathcal{L}^{\text{bce}}+\lambda_{p}\mathcal{L}^{r}$, where $\lambda_{p}$ is a loss weight.
-We employ stage-wise training: first pretraining with only the reconstruction loss ($\mathcal{L}^{r}$), followed by end-to-end training with both losses.
+where $\ell_{\text{bce}}$ is the binary cross-entropy loss function. The total training loss of our framework is $\mathcal{L}=\mathcal{L}^{\text{bce}}+\mathcal{L}^{r}$.
 4 Experiments
 4.1 Datasets
-As no prior dataset exists for the task, we propose two challenging scene datasets, both essential for thorough evaluation: ToysAD-8K and PartsAD-15K.
-ToysAD-8K includes real-world objects from multiple categories.
-This allows us to evaluate our model’s ability to generalize to unseen object categories.
-PartsAD-15K comprises a more diverse collection of mechanical object parts with arbitrary shapes, thus being free from any class-level inductive biases.
+We have proposed two new benchmarks for multi-object AD: ToysAD-8K and PartsAD-15K.
+ToysAD-8K consists of real-world objects from multiple categories that enables to evaluate our model’s ability to generalize to unseen object categories.
+PartsAD-15K comprises a more diverse collection of mechanical object parts with arbitrary shapes, thus being free from any class-level inductive biases (e.g., a sheep with three legs are easy to be recognized as anomaly without comparison) of common objects.
 Both datasets include a wide range of fine-grained anomaly instances motivated by real-world applications in inspection and quality control.
@@
 ToysAD-8K.
-We begin with a subset of $1050$ shapes from the Toys4K dataset [41].
-The subset covers a wide range of objects from $51$ categories.
-We automatically create anomalies for a given 3D shape by applying various deformations to both the geometry and texture.
-This includes generating realistic cracks and fractures using [39], applying random geometric deformations [43] like bumps, bends, and twists, as well as randomly translating, rotating, and swapping materials in different parts of the shapes.
-In total, we generated $2345$ anomaly shapes.
-To generate each scene, we first randomly choose a set of objects consisting of both normal and their anomaly of the same instances.
-We ensure that the majority of objects in each scene are normal.
-We also have few scenes where no anomaly is present.
-The rotational poses for the objects are obtained using rigid body simulation [4].
-The objects are scaled and placed into the scene at random locations, ensuring collisions do not occur.
-We generated total $8K$ scenes.
+We start with a subset of $1050$ shapes from the Toys4K dataset [36], covering a wide range of objects across $51$ categories.
+Anomalies are automatically generated for each 3D shape by applying various deformations to both the geometry and texture.
+These include realistic cracks and fractures using [34], applying random geometric deformations [38] like bumps, bends, and twists, as well as randomly translating, rotating, and swapping materials in different parts of the shapes.
+This process results in a total of $2345$ anomaly shapes.
+To generate each scene, we randomly select a set of objects consisting of both the normal and anomalous variants of the same instance.
+Most scenes contain more normal objects than anomalies and few scenes have no anomalous instances.
+The resting poses for the objects are generated using rigid body simulation [4] by dropping the object to the floor.
+Objects are scaled and placed randomly in the scene while avoiding collisions We generated a total of $8K$ scenes.
 Each scene consists of $3$-$6$ objects rendered in $20$ views.
-For the training set, we randomly select $5K$ scenes from the $39$ categories.
+For the training set, we randomly select $5K$ scenes from $39$ categories.
 We build two disjoint test sets.
 The first one (seen) contains $1K$ scenes from the seen categories but with unseen object instances.
-The second one (unseen) contains $2K$ scenes from rest of the $12$ novel categories.
+The second one (unseen) contains $2K$ scenes from the rest of the $12$ novel categories.
 PartsAD-15K.
 We use a subset of the ABC dataset [23] that consists of $4200$ shapes.
-We follow the strategy above to generate anomalies.
+Anomalies are generated following the same strategy outlined above.
 Additionally, for each shape, we sample geometrically close instances from the dataset and assign them as anomalies to use in the same scene.
-This approach generates a large set of high-quality anomalies that closely resemble their normal counterparts in high-level geometry, but with subtle shape variations that make them anomalous in the context of the normal ones.
-In total, we generated $10,203$ anomaly shapes.
-Using these shapes, we created $15K$ scenes, each consisting of $3$ to $12$ objects rendered from $20$ different viewpoints.
-We divide the dataset into a $12K$ training set and the rest as the test set.
-Note that the proposed datasets, source code, and models will be made public based upon publication.
-4.2 Implementation details
-We use a ResNet50-FPN [30] as our 2D encoder backbone.
-Our 3D backbone consists of a four-scale encoder-decoder-based 3D CNN [32].
-We use $M=5$ images as input, each with a resolution of $256\times 256$.
-Our architecture is flexible to accept a different $M$ during inference.
-During training, we consider a total of $2M$ views, which we separate into two sets, each containing $M$ views.
-We use one set to build the neural volume and the cameras of the other set to render the results, and vice versa.
-The 3D volume $\bm{F}_{v}$ contains $96\times 96\times 16$ voxels with a voxel size of $4cm$.
-We sample $128$ points on each ray for rendering.
-We render the features with a spatial dimension of $32\times 32$.
-The $I_{t\sigma}$ in Eq.
-3 is the ground truth segmentation mask of the input image, we only use it for training.
-The threshold applied to the density volume $\bm{V}_{\sigma}$ is set to be $0.2$ and we run the DBScan algorithm with its default parameters.
-We resize the teacher’s (DINOv2) features to the same spatial dimension for loss computation.
-These features are pre-computed for all scenes using publicly available weights, which are not updated during distillation.
-These DINOv2 features are then reduced to $d_{f}$=$128$ channel dimensions using PCA before distillation.
-We employ three sparse voxel attention blocks, and each applies 8-headed attention.
-The value of $k$ is chosen as $20$.
-The loss weights $\lambda_{\sigma}$, $\lambda_{f}$, and $\lambda_{p}$ are all set to $1$.
-We first pretrain the network with only the reconstruction loss for $50$ epochs.
-Then, we train the network end-to-end with both the reconstruction loss and the binary classification loss for another $50$ epochs.
-We maintain a batch size of $4$, and use the Adam optimizer with a learning rate $2\times 10^{-5}$.
-Finally, the run-time of our method is $65$ms on a single A40 GPU for a typical scene with 5 views as input.
-4.3 Baseline Comparisons
-We quantitatively evaluate the anomaly classification results using two evaluation metrics – the area under the ROC curve (AUC) and accuracy.
-A prediction is considered correct if the bounding box IoU is greater than 0.5 and the corresponding anomaly classification is correct.
-We have not designed a separate localization metric as our estimated bounding boxes are very accurate.
-These metrics are calculated object-wise and then averaged across all test scenes.
-Since no prior work exists on this task, we define several competitive baselines for comparison.
-Specifically, in  Tab.
-1, we compare our method with two relevant approaches: a reconstruction-based baseline and two multi-view 3D object detection methods.
-For the reconstruction-based design, we first employ COLMAP [38] to obtain a point cloud reconstruction of the multi-object scene environment.
-We use default dense reconstruction parameters but utilize the provided ground truth camera matrices.
-Since 5 views are insufficient, we use a total of 20 views to reconstruct the scene.
-We extract individual object point clouds from the reconstructed scenes, and train a Siamese style network to obtain their pairwise similarity.
-We use DGCNN [49] (pretrained on ShapeNet [14]) as the point cloud feature extractor and the triplet loss [22] to supervise the network.
-Finally, we employ a voting strategy to aggregate the pairwise distances of all objects in the scene and obtain their individual binary labels.
-We observe that the accuracy of this method is sensitive to the reconstruction quality, leading to poor performance on both datasets.
-Next, we adapt ImVoxelNet [37] and DETR3D [48], two multi-view object detection frameworks in our problem setting, aiming to locate and classify each object in the scene as either anomaly or normal.
-ImVoxelNet uses the similar 2D-3D projection as ours to construct the voxel representation and then a 3D detection head [29] outputs the final prediction.
-DETR3D is a transformer-based design and uses the set prediction loss [9] for end-to-end detection without NMS.
-We use the ground truth 3D bounding boxes to train these models and the reported scores are calculated object-wise based on the output of the classification head.
-We observe that while both methods perform well for large cracks or fractures, they struggle when intra-group comparison is necessary.
-This is because they tend to memorize certain anomaly types without learning to generalize to compare with other objects in the scene.
-Our method significantly outperforms all baselines (Tab.
-1) on both datasets, highlighting the effectiveness of our dedicated architecture for matching corresponding regions.
-Notably, the performance drop on the unseen set is relatively lower in our case, attributed to the robust 3D representation that effectively generalizes to novel categories.
-We show qualitative results of our method in Fig.
-3
+This results in high-quality anomalies that closely resemble their normal counterparts in overall geometry but differ in subtle shape variations, making them anomalous in the context of the normal ones.
+In total, we generate $10,203$ anomaly shapes and create $15K$ scenes, each consisting of $3$ to $12$ objects rendered from $20$ different viewpoints.
+The dataset is split into a $12K$ training set and the remaining $3K$ scenes (unseen shapes) as the test set.
+Anomaly generation and checking.
+Our anomaly generation process is fully automated with several quality checks to ensure the anomalies are realistic.
+For example, if a part detaches from the main body during positional or rotational deformations, the anomaly is discarded and regenerated with adjusted parameters.
+Similarly, if removing a part makes the shape disconnected (e.g., teddy bear with a floating hand after removing an arm), we discard the anomaly and try a different part.
+For fracture anomalies, if a particular fracture removes more than $90\%$ or less than $10\%$ of an object, it is rejected and regenerated.
+We also ensure that the anomalous region of an object is visible from at least one viewpoint.
+Figure 3: AD qualitative results on the unseen test categories of ToysAD-8K (top row) and the test set of PartsAD-15K (bottom row) using our proposed framework.
+Due to limited space, two views are shown per scene.
+The model prediction is shown with a yellow bounding box: a 3D box for the first example (banana) and a projected box for the others for simplicity.
+Our model successfully predicts the correct object in all cases shown above.
 Table 1: Quantitative results. We compare our method to three related works in two datasets and report the results in terms of anomaly detection AUC and accuracy.
-Datasets COLMAP [38] ImVoxelNet [37] DETR3D [48] Ours
+Datasets Recon-Recog ImVoxelNet [32] DETR3D [43] Ours
 AUC Accuracy AUC Accuracy AUC Accuracy AUC Accuracy
@@
 PartsAD-15K 72.78 61.34 72.80 64.34 74.49 65.11 86.12 79.68
-Figure 3: AD qualitative results on (a) the unseen test categories of ToysAD-8K and (b) the test set of PartsAD-15K using our proposed framework.
-The green box denotes correct predictions, while the red box indicates where the model misses the anomaly.
-Due to limited space, one view is shown per scene.
-Please refer to the supplementary materials for more results.
+4.2 Implementation details
+We use a ResNet50-FPN [25] as the 2D encoder backbone and a four-scale encoder-decoder-based 3D CNN [27] as the 3D backbone.
+For input, we use $M$=$5$ views, each with a resolution of $256\times 256$, though our model can accept a different $M$ during inference.
+During training, we use $2M$ views, divided into two sets of $M$ views.
+One set is used to build the neural feature volume, while the other set’s camera views are used to render the results, and vice versa.
+The 3D volume $\bm{F}_{v}$ contains $96\times 96\times 16$ voxels with a voxel size of $4cm$.
+We sample $128$ points per ray for rendering, and the rendered features ($\hat{\Phi}_{t}$) have a spatial dimension of $32\times 32$ (i.e., $h_{f}\times w_{f}$).
+The ground truth segmentation mask of the input image ($I_{t\sigma}$ in Eq.
+3) is only used during training.
+The threshold applied to the density volume $\bm{V}_{\sigma}$ is set to be $0.2$ and the DBScan algorithm is run with its default parameters.
+We resize the teacher’s (DINOv2) features to the same spatial dimension for loss computation.
+These features are precomputed for all scenes using the publicly available weights, which are not updated during distillation.
+These DINOv2 features are then reduced to $d_{f}$=$128$ channel dimension using PCA before distillation.
+We employ three sparse voxel attention blocks, and each applies 8-headed attention with $k$=$20$.
+The network is first pretrained for $50$ epochs with only the reconstruction loss to build a reliable initialized 3D feature volume, then trained end-to-end with both the reconstruction and the binary classification losses for another $50$ epochs.
+The batch size is $4$, and we use the Adam optimizer with a learning rate $2\times 10^{-5}$.
+The run-time of our method is $65$ms on a single A40 GPU for a typical scene with 5 views as input.
+4.3 Baseline Comparisons
+We evaluate the anomaly classification results using two metrics – the area under the ROC curve (AUC) and accuracy.
+A prediction is considered correct if the bounding box IoU is greater than 0.5 and the anomaly classification is correct.
+We do not use a separate localization metric since our bounding box estimations are highly accurate.
+These metrics are calculated object-wise and then averaged across all test scenes.
+Since no prior work exists for this task, we compare our method with two relevant competitive baselines: a reconstruction-based baseline and two multi-view 3D object detection methods in  Tab.
+1.
+For the reconstruction-based baseline, we first employ COLMAP [33] to generate a point cloud reconstruction of the multi-object scene environment.
+We use default dense reconstruction parameters but utilize the provided ground truth camera matrices.
+Given that 5 views are insufficient, we use a total of 20 views to reconstruct the scene.
+We extract point clouds from the reconstructed scenes for each object, and train a Siamese-style network to obtain their pairwise similarity.
+We use DGCNN [42] (pretrained on ShapeNet [11]) as the point cloud feature extractor and the triplet loss [18] to supervise the network.
+A voting strategy is then applied to aggregate the pairwise distances of all objects in the scene and classify the objects.
+However, this method’s performance is highly sensitive to the reconstruction quality, resulting in poor performance on both datasets.
+Next, we adapt two multi-view object detection frameworks, ImVoxelNet [32] and DETR3D [43], to locate and classify each object in the scene as either anomalous or normal.
+ImVoxelNet constructs a voxel representation using a 2D-3D projection similar to ours, followed by a 3D detection head [24] for the final prediction.
+DETR3D, a transformer-based design, uses the set prediction loss [7] for end-to-end detection without non-maximum suppression.
+Both models are trained on our datasets and their performance is evaluated object-wise based on the output of the classification head, similar to ours.
+After examining the qualitative results (see supplementary), we observe that although both methods perform well for large cracks or fractures, they struggle when intra-group comparison is necessary.
+This is because they tend to memorize certain anomaly types and fail to capture the cross-object relations in the scene.
+Our method significantly outperforms all baselines on both datasets (see Tab.
+1).
+This highlights the effectiveness of our architecture, which is designed to match corresponding regions across objects.
+Notably, our model exhibits a smaller performance drop on the unseen set, thanks to the robust 3D representation that generalizes to novel categories.
+Qualitative results of our method are shown in Fig.
+3
 4.4 Ablations and Model Analysis
-We conduct several studies into the performance of the proposed model including changes to architecture, robustness analysis and real-world experiments.
-Unless stated otherwise, all experiments in this section are carried out on the ToysAD-8K unseen set using 5 input views.
-Figure 4: Correspondences are obtained in 3D space using the neural feature field $\bm{V}_{f}$, then projected onto 2D views using camera matrices for visualization.
-The feature field is rendered at respective viewpoints, the first 3 PCA components are used, with each matched to a different color channel.
+In this section, we present an ablation study of our model, analyze its robustness, and evaluate its performance on real scenes.
+Unless stated otherwise, all experiments are conducted on the ToysAD-8K unseen set using 5 input views.
 Ablation of architecture design.
-We ablate the core components in our model and report the results in Tab.
+We evaluate the main components of our model through an ablation study and report the results in Tab.
 2.
 All variants include the 3D feature fusion module and are optimized at least for the image reconstruction loss, which is essential for constructing scene geometry.
-Variant A directly maps object-centric features to their binary labels using an MLP without comparing them, while for variant B, standard attention layers are applied to the object-centric features to learn cross-object correlations.
+Variant A directly maps object-centric features to their binary labels using an MLP without comparing them.
+In contrast, variant B applies standard attention layers to the object-centric features to learn cross-object correlations.
 Despite the attention layers, B only shows a slight improvement (+2.1% AUC) over A.
-Table 2: Ablation Results on ToysAD-8K
+Table 2: Ablation Results on ToysAD-8K.
 Variants AUC Accuracy
@@
 Final: C + sparse voxel attention 89.15 81.57
-We then introduce DINOv2 feature distillation (variant C), which significantly boosts performance by 5.8% AUC and 11.3% accuracy, indicating the importance of the part-aware 3D representation and correspondences for our task (also see Fig.
-4).
-Our final design achieves the performance gain by utilizing sparse voxel attention, which focuses on the top-$k$ most relevant features.
+We then introduce DINOv2 feature distillation (variant C), which significantly improves performance by 5.8% AUC and 11.3% accuracy.
+This demonstrates the importance of the part-aware 3D representation and correspondences for our task (see Fig.
+6).
+Our final design achieves further performance gain by using sparse voxel attention, which focuses on the top-$k$ most relevant features.
 This leverages robust correspondences learned through DINOv2 feature distillation, effectively eliminating noisy correlations and directing attention solely to corresponding object regions.
 Robustness.
-Our method achieves some robustness to occlusion by effectively using input from multiple viewpoints.
-Fig.
-7 illustrates examples where our model accurately identifies the anomalous region, even when occluded in some views.
+Our method demonstrates robustness to occlusions by effectively utilizing input from multiple viewpoints.
+As shown in Fig.
+4, our model accurately identifies the anomalous region even when parts of the object are occluded in some views.
 In Fig.
-5 (left), we show how well our model can perform with additional or reduced input views at test time.
-We train our model with 5 views and then evaluate it using 1, 3, 5, 10 and 20 views.
-These results clearly indicate that our model can perform reasonably well with only 3-5 views; however, additional views can be used to boost performance at test time.
-Figure 5: Impact of number of views (left) and object count (right) on model performance.
+5 (left), we show how well our model can perform with varying numbers of input views during testing.
+Trained with 5 views, the model is evaluated using 1, 3, 5, 10, and 20 views.
+The results clearly show that our model performs well with only 3-5 views, while additional views further enhance the performance.
+Figure 4: Resolving occlusion and 3D ambiguity using multi-view images.
+The anomaly ‘sheep’ in top has a missing tail (only visible in the 2nd view due to occlusion), and the ‘hammer’ handle in bottom is bent (only apparent from the 2nd view-angle due to 3D ambiguities).
+Figure 5: Impact of the number of views (left) and object count (right) on model performance.
+Figure 6: Correspondences are obtained in 3D using the neural feature field $\bm{V}_{f}$ and projected onto 2D views via camera matrices for visualization.
+The feature field is rendered at respective viewpoints, with the first 3 PCA components mapped to different color channels.
 We analyze the impact of object count on model performance (see Fig.
 5 (right)) using the PartsAD-15K dataset.
-Increasing the number of objects in a scene usually leads to more occlusion and lower resolution for each individual object.
+As the number of objects in a scene increases, occlusion and lower resolution for individual objects typically occur.
 However, the performance drop between the two extremes is minimal ($<$ 1.5% AUC) as shown in the figure.
-In addition, we investigate the model’s ability to generalize to more objects at test time than it observed during training.
+We also assess the model’s ability to generalize to scenes with more objects than it encountered during training.
 To this end, we trained our model on scenes with 3-7 objects and tested on two sets: one with 3-7 objects (AUC: 86.75) and another with 8-12 objects (AUC: 85.10).
 These results demonstrate our model’s adaptability to varying object counts.
-In another study, we test our primary objective of choosing the majority group as normal and the rest as anomalies using an example shown in Fig.
-6.
-We create five scenes (a-e) by using two geometrically similar objects, and gradually introduce more from one (1 to 5 respectively) while maintaining a total of six objects.
-For example, in scene (a), the first object appears only once, making it an anomaly.
-Similarly, in scene (e), the second object appears only once, hence considered an anomaly.
+In another experiment, we assess our model’s performances when varying the ratio of normal to anomalous instances during testing as shown in Fig.
+8.
+We create five scenes (a-e) by using two geometrically similar objects ob1 and ob2, and gradually introduce more instances of ob1 (1 to 5 respectively) while maintaining a total of six objects.
+For example, in scene (a), ob1 appears only once, making it an anomaly.
+Similarly, in scene (e), ob2 appears only once, hence considered an anomaly.
 We utilize a consistent background across all scenes to ensure uniformity.
@@
 Real world testing.
-Here we apply our model, which is trained on the synthetic dataset, on a small set of real test scenes.
-Each scene is set up in an indoor environment with adequate lighting and is captured using a 3D scanning software [1].
-This results in a set of input views with globally optimized cameras.
+We apply our model, trained on the synthetic dataset, to a small set of real test scenes (a tabletop with three cups, four coca-cola cans, four medicine bottles, or three bananas).
+Each scene is captured in an indoor environment with adequate lighting using a 3D scanning software [1].
+We obtain a set of input views of the scenes with globally optimized camera intrinsic and extrinsic parameters, which are then fed into our trained model.
+The results are shown in Fig.
+7.
+Our model is able to correctly distinguish the anomaly objects in all four scenes.
+Figure 7: Four real-world scenes are tested using our method and it can successfully detect all anomalous instances. Five views are used for testing, but only one view per scene is shown.
+Limitations.
+Our benchmark and model have a few limitations.
+First, this work focuses on a specific set of anomalies common in manufacturing scenarios, potentially overlooking other types of anomalies found in different applications.
+Due to the high cost and difficulty of obtaining real damaged objects, our dataset primarily uses shapes of synthetic objects.
+Additionally, our model assumes that object instances are rigid and does not handle articulations or deformations.
+It also assumes objects are not neither touching nor fully occluded in the scene.
+Furthermore, when a scene contains mostly anomalies, the lack of ‘normal’ data points for comparison (except in cases like fractures or cracks) can make detection more challenging.
+Lastly, during real-world scene testing, factors such as noisy camera poses and the gap between synthetic and real environments may degrade the performance.
 Fig.
-8 illustrates the results for three such scenes.
-Limitations.
-Our benchmark and model also have a few limitations.
-Firstly, this work focuses solely on a limited set of anomalies that are common to manufacturing scenarios; potentially missing other anomalies in real-world scenarios.
-Since acquiring real damaged objects is expensive and difficult, our dataset primarily uses shapes of synthetic objects.
-Moreover, our model assumes that object instances are rigid and cannot handle articulations or deformations.
-It also assumes that objects are not touching or completely occluded in space.
-Moreover, a scene with mostly anomalies can be challenging for our model due to the lack of ‘normal’ data points for comparison (except for fractures or cracks).
-Moreover, the performance is dependent on the anomalous region being captured in at least one view.
-Finally, in real-world scene testing, the noisy camera poses as well as the gap between synthetic and real-world environments may degrade the performance.
+9 shows some failure cases of our model.
 Figure 8: We create five scenes (a-e) with two similar objects, varying their counts per scene while maintaining a constant total.
-Our model correctly identifies minorities as anomalies in all cases except (c), where the equal number of objects creates ambiguity.
+Our model correctly identifies minorities as anomalies in all cases except (c), where an equal number of objects creates ambiguity.
 Our model still selects one group.
-Figure 7: Resolving (a) occlusion and (b) 3D ambiguity using multi-view images.
-The anomaly ‘sheep’ in (a) has a missing tail (only visible in the last view due to occlusion), and the ‘hammer’ handle in (b) is bent (only apparent from the last view-angle due to 3D ambiguities).
-Figure 8: Three real-world scenes are tested using our method and it can successfully detect all anomalous instances.
-Five input views are used for testing, while one view for each scene is used for visualization.
+The yellow box indicates the predicted anomaly object.
+Figure 9: Failure cases.
+We use yellow box for our model’s prediction and red box for the ground truth.
+In the first example, our model incorrectly predicts an anomaly, while in the second, it correctly identifies two objects but misses the third anomaly.
 5 Conclusion
 In this paper, we have introduced a novel AD problem inspired by real-world applications along with two new benchmarks.
-The proposed task goes beyond the traditional AD setting and involves a cross study of objects in a scene from multiple camera viewpoints to identify the ‘odd-looking’ minority group.
+The proposed task goes beyond the traditional AD setting and involves a cross study of objects in a scene from multiple camera viewpoints to identify the ‘odd-looking’ instances.
 We show that our model is robust to varying number of views and objects, and outperforms the baselines that do not consider cross-object correlations.
-Broader Impacts.
-The proposed techniques could be potentially used in improving quality checks and product safety by providing warnings to the users in production lines.
-The authors are not aware of any potential harm that may arise when the technology is used.
-Appendix A Appendix / supplemental material
-This appendix is structured as follows: we present additional details of the proposed datasets in Sec. A.1, training details in Sec. A.2, and additional qualitative results in Sec. A.3.
-A.1 Data Generation Details
-Our proposed scene AD datasets, ToysAD-8K and PartsAD-15K, are built upon two publicly available 3D shape datasets: Toys4K [41] (Creative Commons and royalty-free licenses) and ABC [27] (MIT license).
-For the ToysAD-8K, we selected 1,050 shapes from the Toys4K dataset, focusing on the most common real-world objects across 51 categories.
-A complete list of these categories is provided in Table 3.
-On the other hand, PartsAD-15K is a non-categorical dataset.
-For this dataset, we randomly selected a subset of 4,200 shapes from the large-scale ABC.
-For both datasets, we consider the following anomaly types: cracks, fractures, geometric deformations (e.g., bumps, bends, and twists), translation, rotation, material mismatch, and missing parts.
-Additionally, for PartsAD-15K, we use a different but geometrically similar instance in the same scene as an anomaly.
-To obtain a geometrically similar shape, we use feature-based KNN clustering.
-Specifically, we extract DINOv2 features from multi-view images (rendered from multiple fixed viewpoints) of the 3D shapes, concatenate these features, and use the resulting concatenated features to build the KNN cluster.
-Then, we retrieve a similar 3D shape by querying the KNN cluster.
-To ensure the retrieved shape is geometrically similar, we calculate the Chamfer distance between the shapes and only accept a shape if the distance is below a certain threshold.
-Table 3: Dataset composition of ToysAD-8K
-Categories
-Seen dinosaur, fish, frog, monkey, light, lizard, orange, boat, dog, lion, pig, cookie, panda, chicken,
-orange, ice, horse, car, airplane, cake, shark, donut, hat, cow, apple, bowl, hamburger, octopus,
-giraffe, chess, bread, butterfly, cupcake, bunny, elephant, fox, deer, bus, bottle
-Unseen mug, plate, robot, glass, sheep, shoe, train, banana, cup, key, penguin, hammer
-Our anomaly generation process is automatic.
-To ensure the quality of the generated anomalies, we perform several checks.
-For example, during positional or rotational anomaly creation, if a part detaches from the main body during deformation, we reject the anomaly and try again with adjusted parameters.
-Similarly, if removing a part makes the shape impractical, we discard it and try removing a different part.
-For fracture anomaly, if a particular fracture removes more than 90% or less than 10% of an object we discard the sample and regenerate another fracture.
-Finally, we ensure the anomalous region of an object is visible from at least one viewpoint.
-To generate a realistic scene environment, we use PBR materials [44] for floors and HDRI environment maps [56] for image-based lighting to illuminate scenes.
-We randomly select a pair of PBR material and HDRI environment maps from the assets to randomize the scene background.
-Objects are placed randomly such that no collisions occur, and each object’s rotational poses are obtained using Blender’s rigid body simulation [43].
-We employed Blender 2.93 [43] with Cycles ray-tracing renderer for photo-realistic rendering.
-Blender 2.93 is released under the GNU General Public License (GPL, or “free software”), and the PBR and HDRI maps are released under the CC0 license.
-Our framework can easily be trained with real-world manufacturing scene environments.
-Our model relies solely on 2D supervision, making the data collection process much easier.
-We also do not need precise annotation of 3D bounding boxes as they are not used during training.
-For annotating instance-wise anomaly labels, we can use 2D bounding boxes, which can be projected in 3D using Visual Hull [28], then used for coarse localization.
-A.2 Training Details
-We train our model in two stages.
-In the first stage, we train it with just image and feature reconstruction losses.
-In the second stage, we train the model end-to-end with both reconstruction and binary classification losses.
-All experiments are performed on a single NVIDIA A40 GPU with a batch size of 4, utilizing 28GB of GPU memory.
-The first stage takes 36 hours to complete, followed by an additional 24 hours for the second stage.
-The ablation experiments are conducted on the same workstation with the same GPU by removing one or a few core components from the full method.
-Specifically, variant methods A and B take $36$ hours to train the first stage, while only taking $14$ hours to train the second stage.
-Regarding variant method C, it takes similar $36$ and $24$ hours for the two stages as in the full method.
-We compare our method with COLMAP (BSD license), ImVoxelNet (MIT license), and DETR3D (MIT license).
-For the COLMAP-based approach, we use DGCNN (MIT license) as a feature extractor.
-In terms of training time of these baseline methods, COLMAP takes $10$ hours to train, ImVoxelNet takes $24$ hours to train, and DETR3D takes $2$ days to converge.
-The standard deviation of all our experiments (including the ablations and our method) under multiple runs is less than 0.5.
-A.3 Additional Qualitative Results
-In Fig. 9 and Fig. 10, we present additional qualitative results on ToysAD-8K and PartsAD-15K, respectively.
-Figure 9: Additional results on the unseen set of ToysAD-8K dataset.
-For rows 1 to 3, the anomalies are easy to spot and self-explanatory.
-In row 4, there are two anomalies: one with broken outer parts at the back (see view 5), and another with a tilted roof (see views 1 and 2).
-For row 5, one leg is broken (see view 2).
-In row 6, the eye is missing (see view 5).
-Figure 10: Additional results on the PartsAD-15K dataset.

### v4

@@
 Abstract
-This paper introduces a novel anomaly detection (AD) problem that focuses on identifying ‘odd-looking’ objects relative to the other instances in a given scene.
-In contrast to the traditional AD benchmarks, anomalies in our task are scene-specific, defined by the regular instances that make up the majority.
-Since object instances may be only partly visible from a single viewpoint, our setting employs multiple views of each scene as input.
-To provide a testbed for future research in this task, we introduce two benchmarks, ToysAD-8K and PartsAD-15K.
-We propose a novel method that constructs 3D object-centric representations from multiple 2D views for each instance and detects the anomalous ones through a cross-instance comparison.
-We rigorously analyze our method quantitatively and qualitatively on the presented benchmarks.
-The datasets, source code, and models will be made available upon publication.
+This paper introduces a novel anomaly detection (AD) problem aimed at identifying ‘odd-looking’ objects within a scene by comparing them to other objects present.
+Unlike traditional AD benchmarks with fixed anomaly criteria, our task detects anomalies specific to each scene by inferring a reference group of regular objects.
+To address occlusions, we use multiple views of each scene as input, construct 3D object-centric models for each instance from 2D views, enhancing these models with geometrically consistent part-aware representations.
+Anomalous objects are then detected through cross-instance comparison.
+We also introduce two new benchmarks, ToysAD-8K and PartsAD-15K as testbeds for future research in this task.
+We provide a comprehensive analysis of our method quantitatively and qualitatively on these benchmarks.
 Figure 1: (a) We propose a new anomaly detection task focused on identifying ‘odd-looking’ objects relative to other instances within a scene.
@@
 1 Introduction
-Anomaly detection (AD) [10, 29] aims to detect patterns that deviate from expected behavior.
-In standard AD benchmark for computer vision, anomalies may arise from high-level factors such as introduction of an object instance from an unseen category [9, 31, 2] or low-level ones such as defects and variations in the normal object shape and texture [8, 6, 13].
-In these benchmarks, AD methods analyze each object individually, ignoring relationships between objects in the same scene or contextual factors.
-However, in many real-world scenarios, ‘normality’ can be conditional.
+Anomaly detection (AD) [11, 31] identifies patterns deviating from expected behavior.
+In standard vision AD benchmarks, anomalies arise from high-level factors such as presence of a previously unseen category [10, 33, 2] or low-level ones such as defects and variations in the normal object shape and texture [9, 6, 14].
+These benchmarks typically assess objects in isolation, disregarding inter-object relationships and contextual dependencies.
+However, in many real-world scenarios, ‘normality’ is often context-dependent.
 For instance, in a production line for blue-handled coffee cups, a yellow handle is anomalous, and in a production line for yellow-handled coffee cups, a blue handle is anomalous (see Fig.
 1(b)).
-This context dependent normality cannot be addressed by methods assuming a fixed standard of normality.
-Inspired by real-world visual inspection, we introduce a new AD problem with two new benchmarks and a novel solution.
-In our setup (see Fig.
-1(a)), each image contains a scene with multiple instances of the same object (e.g., coffee cup) including either all normal, or a mix of normal and anomalous instances.
-Since defining exact shape and appearance specifications on paper is often impractical, we assume each scene has multiple normal instances together providing a scene-specific reference for ‘normality’.
-Our objective is to detect anomalous instances in a scene and to generalize to previously unseen scenes including unseen objects and layouts.
-Unlike existing benchmarks, this task often requires a holistic understanding of the scene by comparing instances with each other with the exception of few anomalies such as cracks and misaligned parts that can be identified by inspecting individual instances.
-Additionally, since performing AD from a single view can be ambiguous due to self-occlusion and occlusion between the instances, we provide multiple views of the scene to cover its entire relevant extent unlike existing benchmarks with single-view inputs (see Fig.
+Methods assuming a fixed standard of normality cannot accommodate such context-dependent anomalies.
+Inspired by real-world visual inspection, where efficiency in automatically processing large volumes of the same product is crucial, we introduce a new AD problem with along two new benchmarks and a novel solution.
+Our setup (see Fig.
+1(a)) consists of scenes containing multiple instances of the same object (e.g., coffee cup) including either all normal, or a mix of normal and anomalous instances.
+Each scene has at least two normal instances, which together serve as a scene-specific reference for ‘normality’.
+The objective is to detect anomalous instances while generalizing to previously unseen scenes including unseen objects and layouts.
+This approach facilitates cost-effective models that do not require retraining for new products.
+Our task diverges from traditional AD tasks in two key ways: (1) anomaly detection is scene-specific, requiring a comparative analysis of instances rather than solely relying on individual inspection – only certain anomalies (e.g., cracks and misaligned parts) are detectable at instance-level; (2) it involves multi-view scene analysis, capturing a comprehensive object model while addressing self-occlusion and inter-instance occlusion, unlike standard AD benchmarks that typically rely on single-view inputs (see Fig.
 1(a)).
-The proposed problem presents several challenges requiring the following key capabilities: i) achieving 3D understanding of the scene and view registration from multiple camera viewpoints without groundtruth 3D knowledge while handling occlusions, ii) aligning and comparing object instances without their pose information during training and testing, iii) learning representations that generalize to unseen object instances from seen or unseen categories at test time.
-To tackle these challenges, we introduce a novel method that takes as input multiple views of the same scene, projects them into a 3D voxel grid, generates a 3D object-centric representation for each instance, and predicts their labels by cross-correlating instances using an efficient attention mechanism.
-We utilize advances in differentiable rendering [26] and self-supervised learning [28] to supervise 3D representation learning.
-Specifically, we render the voxel representations for multiple viewpoints and match them with the given 2D views to produce geometrically consistent features.
-Additionally, we encourage our model to learn part-aware 3D representations by distilling features from 2D self-supervised model DINOv2 [28], enhancing the correspondence matching across instances and their alignment.
-Finally, since there is no prior benchmark for this task, we propose two new benchmarks – ToysAD-8K and PartsAD-15K, which include common objects categories as well as mechanical parts respectively to facilitate further research.
-Our method significantly outperforms baselines that do not compare instances with each other, and we rigorously analyze various aspects of our model and benchmarks.
+The proposed problem presents several unique challenges requiring the following key capabilities: i) 3D-aware modeling: constructing accurate representations of object instances by capturing their geometry and appearance from multiple viewpoints while handling occlusions, ii) pose-agnostic comparison: aligning and comparing object instances effectively without their pose information during training and testing, iii) generalizable representations: learning representations that generalize to unseen object instances from both seen or unseen categories at test time.
+To tackle these challenges, we introduce a novel method that integrates multi-view scene analysis by projecting them into a 3D voxel grid and generating 3D object-centric representations.
+Our model classifies instances through cross-instance correlation using an efficient attention mechanism.
+Leveraging recent advances in differentiable rendering [28] and self-supervised learning [30], we ensure robust 3D representation learning by rendering voxel representations for multiple viewpoints and aligning them with corresponding 2D views for geometric consistency.
+Additionally, we enhance part-aware 3D representation learning by distilling features from 2D self-supervised model DINOv2 [30], improving correspondence matching and pose-invariant alignment.
+As no existing benchmarks address this task, we introduce two new benchmarks: ToysAD-8K including common objects categories, and PartsAD-15K focusing on mechanical parts, to foster further research.
+Our method significantly surpasses baselines that lack cross-instance comparison and part-awareness, and we provide an in-depth analysis of our model.
 2 Related Work
 AD benchmarks.
-A key challenge in AD research is the scarcity of large datasets containing realistic anomalies.
+A major challenge in AD research is the limited availability of large datasets with realistic anomalies.
 Earlier works [10, 33] focusing on high-level semantic anomalies often use existing classification datasets by treating a subset of classes as anomalies and the remainder as normal.
-There also exist several datasets containing real-world anomaly instances.
-For example, MVTec-AD [5] includes industrial objects with various defects like scratches, dents, and contaminations, Carrera et al.
-missing[8] presents various defects in nanofibrous material, and VisA [52] comprises complex industrial objects such as PCBs, as well as simpler objects like capsules and cashews, spanning a total of 12 categories.
-These datasets assume that objects are pose-aligned, and both normal images and their anomaly counterparts have the same pose.
+There are also datasets containing real-world anomaly instances; for instance, MVTec-AD [5] includes industrial objects with various defects like scratches, dents, and contaminations, while Carrera et al.
+missing[9] presents various defects in nanofibrous material.
+The VisA [56] comprises complex industrial objects such as PCBs, as well as simpler objects like capsules and cashews, spanning a total of 12 categories.
 Zhou et al.
-missing[51] propose a pose-agnostic framework by introducing the PAD dataset, which comprises images of 20 LEGO bricks of animal toys from diverse viewpoints/poses.
-Unlike the works discussed above, we focus on AD in multi-object multi-view scene environments, where anomalies are predicted by assessing their mutual similarity with other objects in the scene.
-The proposed setting enables our framework to work seamlessly on novel object instances without requiring further training.
+missing[55] introduced a pose-agnostic framework by introducing the PAD dataset, which comprises images of 20 LEGO bricks of animal toys from diverse viewpoints/poses.
+Bhunia et al.
+missing[7] proposed a conditional AD detection technique by comparing a given 3D reference shape to a single object in a query image.
+In contrast, we target AD within multi-object, multi-view scene environments, where anomalies are defined relative to scenes and detected through cross-instance in the scene.
+Moreover, our method does not require 3D reference models which can be costly and tedious to obtain.
 Few-shot AD.
@@
 $\psi:\{(\bm{I}_{t},\bm{P}_{t})\}_{t=1}^{M}\mapsto\{({y}_{n},\bm{b}_{n})\}_{n=1}^{N}.$ (1)
-Note that labels $y_{n}$ are defined relative to other objects in the scene.
-For example, consider a group of three coffee cups, of which two have yellow handles, and one has a blue handle.
-In this example, the latter cup is considered an anomaly.
+Importantly $y_{n}$ is defined relative to other objects in the scene.
+For example, consider a group of three coffee cups, where two have yellow handles, and one has a blue handle.
+Here, the blue-handled cup is classified as an anomaly.
+Notably, we do not assume that normal instances outnumber anomalies; rather, we assume that anomalies exhibit instance-specific variations (e.g.,, fractures differing in location and lengths), while normal instances maintain high visual similarity.
 Our pipeline, illustrated in Fig.
@@
 4.1 Datasets
-We have proposed two new benchmarks for multi-object AD: ToysAD-8K and PartsAD-15K.
-ToysAD-8K consists of real-world objects from multiple categories that enables to evaluate our model’s ability to generalize to unseen object categories.
-PartsAD-15K comprises a more diverse collection of mechanical object parts with arbitrary shapes, thus being free from any class-level inductive biases (e.g., a sheep with three legs are easy to be recognized as anomaly without comparison) of common objects.
-Both datasets include a wide range of fine-grained anomaly instances motivated by real-world applications in inspection and quality control.
-Scenes are generated with diverse backgrounds, illuminations, and camera viewpoints using photo-realistic ray tracing [38].
+ToysAD-8K comprises real-world objects from multiple categories, enabling evaluation of our model’s generalization to unseen object categories.
+PartsAD-15K includes a diverse collection of mechanical parts with arbitrary shapes, eliminating class-level inductive biases present in common objects (e.g., a sheep with three legs are easy to be recognized as anomaly without comparison).
+Both datasets feature fine-grained anomaly instances inspired by real-world applications in inspection and quality control applications.
+Scenes are rendered with diverse backgrounds, illuminations, and camera viewpoints using photo-realistic ray tracing [40].
 Next, we discuss the data generation for both datasets.
@@
 This process results in a total of $2345$ anomaly shapes.
-To generate each scene, we randomly select a set of objects consisting of both the normal and anomalous variants of the same instance.
-Most scenes contain more normal objects than anomalies and few scenes have no anomalous instances.
-The resting poses for the objects are generated using rigid body simulation [4] by dropping the object to the floor.
+Each scene is generated by randomly selecting a set of objects, including both the normal and anomalous variants of the same instance.
+While most scenes contain more normal objects than anomalies, some scenes includes only normal instances.
+Object resting poses are determined through rigid body simulation [4] where objects are dropped onto the floor.
 Objects are scaled and placed randomly in the scene while avoiding collisions We generated a total of $8K$ scenes.
@@
 Limitations.
-Our benchmark and model have a few limitations.
-First, this work focuses on a specific set of anomalies common in manufacturing scenarios, potentially overlooking other types of anomalies found in different applications.
-Due to the high cost and difficulty of obtaining real damaged objects, our dataset primarily uses shapes of synthetic objects.
-Additionally, our model assumes that object instances are rigid and does not handle articulations or deformations.
-It also assumes objects are not neither touching nor fully occluded in the scene.
-Furthermore, when a scene contains mostly anomalies, the lack of ‘normal’ data points for comparison (except in cases like fractures or cracks) can make detection more challenging.
-Lastly, during real-world scene testing, factors such as noisy camera poses and the gap between synthetic and real environments may degrade the performance.
+Our benchmark and model come with a few limitations.
+First, this work targets specific anomalies common in manufacturing, which may limit its applicability to other types of anomalies found in different settings.
+Given the challenges and cost of obtaining real damaged objects, our dataset primarily uses shapes of synthetic objects.
+Additionally, our model assumes rigid objects and does not account for articulations or deformations.
+It also relies on objects being not touching nor not fully occluded in the scene.
+Furthermore, when a scene contains mostly anomalies, the lack of regular instances for comparison (except in cases like fractures or cracks) can further complicate the detection.
+Finally, real-world scene testing may introduce factors like noisy camera poses and the gap between synthetic and real environments, which could impact the performance.
 Fig.
-9 shows some failure cases of our model.
+9 illustrates some failure cases of our model.
+Most failure cases stem from obscure anomalies due to factors like low resolution, poor lighting, unresolved occlusions in cluttered scenes, limited multi-view images, or challenging objects (e.g., transparent textures).
 Figure 8: We create five scenes (a-e) with two similar objects, varying their counts per scene while maintaining a constant total.
@@
 5 Conclusion
-In this paper, we have introduced a novel AD problem inspired by real-world applications along with two new benchmarks.
-The proposed task goes beyond the traditional AD setting and involves a cross study of objects in a scene from multiple camera viewpoints to identify the ‘odd-looking’ instances.
-We show that our model is robust to varying number of views and objects, and outperforms the baselines that do not consider cross-object correlations.
+We have introduced a novel AD problem inspired by real-world applications along with two new benchmarks.
+The proposed task goes beyond the traditional AD setting and involves a cross study of objects in a scene from multiple viewpoints to identify the ‘odd-looking’ instances.
+We show that our model is robust to varying number of views and objects, and outperforms the baselines that do not consider cross-object correlations and part-aware representations.
+Acknowledgment. The authors acknowledge the support of Toyota Motor Europe.
+Supplementary Material
+This supplementary is structured as follows: we present additional details of the proposed datasets in  Appendix A, training details in  Appendix B, and additional results in  Appendix C.
+Appendix A Data Generation Details
+Our proposed scene AD datasets, ToysAD-8K and PartsAD-15K, are built upon two publicly available 3D shape datasets: Toys4K [38] (Creative Commons and royalty-free licenses) and ABC [24] (MIT license).
+For the ToysAD-8K, we selected 1,050 shapes from the Toys4K dataset, focusing on the most common real-world objects across 51 categories.
+A complete list of these categories is provided in Fig.
+A2, and the split of seen/unseen categories is shown in Tab.
+A1.
+On the other hand, PartsAD-15K is a non-categorical dataset.
+For this dataset, we randomly selected a subset of 4,200 shapes from the large-scale ABC.
+The test set of PartsAD-15K is comprised of unseen shapes.
+Given a 3D mesh model of an object, we automatically create anomalies by applying geometric deformations as described below.
+Table A1: Dataset composition of ToysAD-8K
+Categories
+Seen dinosaur, fish, frog, monkey, light, lizard, orange, boat, dog, lion, pig, cookie, panda, chicken,
+orange, ice, horse, car, airplane, cake, shark, donut, hat, cow, apple, bowl, hamburger, octopus,
+giraffe, chess, bread, butterfly, cupcake, bunny, elephant, fox, deer, bus, bottle
+Unseen mug, plate, robot, glass, sheep, shoe, train, banana, cup, key, penguin, hammer
+Figure A1: Distribution of anomaly types in the proposed ToysAD-8K and PartsAD-15 datasets.
+Figure A2: Distribution of object categories in ToysAD-8K dataset.
+Figure A3: Anomaly detection performance (Accuracy) by anomaly type on the ToysAD-8K and PartsAD-15 datasets.
+Different types of Anomalies.
+We consider the following anomaly types: cracks, fractures, geometric deformations (e.g., bumps, bends, and twists), misaligned parts (e.g., translation and rotation), material mismatch, and missing parts.
+Fig.
+A1 depicts the distribution of different anomaly types in our proposed dataset.
+Before applying deformations, we normalize the mesh vertices to $[-1,1]$.
+Then we synthetically generate realistic cracks and fractures following [36] to mimic the shape’s most geometrically natural breaking patterns.
+For geometric deformations, we use Blender’s [40] Simple Deform and Hook modifier to create global (e.g., bending and twisting) and local deformations (e.g., surface bumps), respectively.
+For translation, we randomly translate the vertices of a random part by an offset.
+The translation offset for each axis is randomly sampled from a uniform distribution of range $[0.04,0.08]$.
+For rotational anomalies, we apply a 3D rotational transformation to a random part.
+The rotation matrix is formed using a random rotation axis and a radian angle sampled from a uniform distribution of range $[0.2,0.4]$.
+The center of rotation is set to a fixed point at one of the connecting points between the anomalous part and the main body of the object.
+We also randomly change the texture of a part or region to create a material mismatch anomaly or completely remove a part to simulate a missing parts anomaly.
+Additionally, for PartsAD-15K, we use a different but geometrically similar instance in the same scene as an anomaly (referred to as ‘distinct’ in Fig.
+A1).
+To retrieve a geometrically similar shape, we use feature-based KNN clustering.
+Specifically, we extract DINOv2 features from multi-view images (rendered from multiple fixed viewpoints) of the 3D shapes, concatenate these features, and use the resulting concatenated features to build the KNN cluster.
+Then, we retrieve a similar 3D shape by querying the KNN cluster.
+To ensure the retrieved shape is geometrically similar, we calculate the Chamfer distance between the shapes and only accept a shape if the distance is below a certain threshold.
+Scene construction.
+For generating scenes, we use resulting anomalies along with their normal instances.
+A scene can have more than one anomaly of the same instance, or it may not have any anomaly at all.
+For randomized object placement, we first determine each object’s resting pose by simulating a rigid body drop using Blender [40].
+This simulation process is repeated for each object instance in the scene, including both normal and anomaly instances.
+Then, we place each object instance randomly using its obtained resting pose, while also ensuring that they do not collide with each other.
+We prevent collisions by maintaining a minimum distance between each pair of objects.
+Quality checks.
+Our anomaly generation process is fully automated, with multiple quality checks to ensure reliable samples.
+For positional or rotational anomaly creation, we discard and regenerate any instance where a part detaches from the main body during deformation.
+Similarly, if removing a part results in an impractical shape, we try a different part instead.
+For fracture anomalies, we discard the sample where the fracture removes more than 90% or less than 10% of an object, regenerating a new fracture in these cases.
+Finally, we ensure the anomalous region of an object is visible from at least one viewpoint.
+Assets.
+To generate a realistic scene environment, we use PBR materials [41] for floors and HDRI environment maps [53] for image-based lighting to illuminate scenes.
+We randomly select a pair of PBR material and HDRI environment maps from the assets to randomize the scene background.
+We employed Blender 2.93 [40] with Cycles ray-tracing renderer for photo-realistic rendering.
+Blender 2.93 is released under the GNU General Public License (GPL, or “free software”), and the PBR and HDRI maps are released under the CC0 license.
+Rendering and view selection.
+We render each scene from various viewpoints sampled from a hemisphere around the scene origin.
+The viewpoint is parameterized in spherical coordinates where azimuth values are sampled uniformly over [0, $2\pi$) and elevation values are uniformly sampled in [$\pi/9$, $92\pi/$ 9].
+We adjust the radius in the range [1.5, 2.5] to ensure that all the objects in the scene are visible (even if partially) in the rendered view.
+Our framework can easily be trained with real-world manufacturing scene environments.
+Our model relies solely on 2D supervision, making the data collection process much easier.
+We also do not need precise annotation of 3D bounding boxes as they are not used during training.
+For annotating instance-wise anomaly labels, we can use 2D bounding boxes, which can be projected in 3D using Visual Hull [25], then used for coarse localization.
+Appendix B Training Details
+We train our model in two stages.
+In the first stage, we train it with just image and feature reconstruction losses.
+In the second stage, we train the model end-to-end with both reconstruction and binary classification losses.
+All the experiments are performed on a single NVIDIA A40 GPU with a batch size of 4, utilizing 28GB of GPU memory.
+The first stage takes 36 hours to complete, followed by an additional 24 hours for the second stage.
+The ablation experiments are conducted on the same workstation with the same GPU by removing one or a few core components from the full method.
+Specifically, variant methods A and B take $36$ hours to train the first stage, while only taking $14$ hours to train the second stage.
+Regarding variant method C, it takes similar $36$ and $24$ hours for the two stages as in the full method.
+We compare our method with Recons-Recog [35], ImVoxelNet [34], and DETR3D [46].
+For the Recons-Recog approach, we use DGCNN [45] as a feature extractor.
+In terms of training time of these baseline methods, Recons-Recog takes $10$ hours to train, ImVoxelNet takes $24$ hours to train, and DETR3D takes $2$ days to converge.
+Figure A4: We qualitatively compare our method with two baselines on the PartsAD-15K dataset.
+ImVoxelNet failed to detect the anomalous object in the first example.
+We use the red box for wrong predictions and the green box for correct predictions.
+Appendix C Additional Results
+Fig.
+A4 shows a qualitative comparison with two baseline methods: ImVoxelNet and DETR3D.
+Our method accurately predicts the anomalies, while the baselines perform poorly.
+In the first three examples, the baselines fail due to their inability to compare with other objects in the scene, which is necessary for correct predictions.
+In the last example, both DETR3D and our method predict correctly, while ImVoxelNet fails.
+In Fig.
+A5 and Fig.
+A6, we present additional qualitative results on ToysAD-8K and PartsAD-15K, respectively.
+We show a breakdown of accuracy across different anomaly types in Fig.
+A3 to interpret the model’s performance in various such scenarios.
+Figure A5: Additional results on the unseen set of ToysAD-8K dataset.
+Each row shows multiple views of the input scene and the green box denotes our model’s prediction.
+For rows 1 to 3, the anomalies are easy to spot and self-explanatory.
+In row 4, there are two anomalies: one with broken outer parts at the back (see view 5), and another with a tilted roof (see views 1 and 2).
+For row 5, one leg is broken (see view 2).
+In row 6, the eye is missing (see view 5).
+Figure A6: Additional results on the PartsAD-15K dataset. Each row shows multiple views of the input scene and the green box denotes our model’s prediction.